### v1

# Kazhdan groups of dimension $16$ with an uncountable range of second $\ell^{2}$-Betti numbers

Abstract
We construct a family of simple, lacunary hyperbolic groups with property $(T)$ that have rational cohomological dimension $16$ and uncountably many values of second $\ell^{2}$-Betti numbers. Along the way, we present new constructions of measurably diverse finitely generated groups, and we prove that the second $\ell^{2}$-Betti number is far from being semi-continuous in the space of marked groups, even assuming good finiteness properties.
1 Introduction
In relation to the Atiyah Conjecture [Ati76], there have been many constructions of matrices over group rings of finitely generated groups whose kernels have irrational von Neumann dimension [Aus13, Gra14, Gra16]. Much less is known about $\ell^{2}$-Betti numbers of groups: the main open question is whether there exists a group of type $FP_{n+1}(\mathbb{Q})$ with irrational $n$-th $\ell^{2}$-Betti number [Gra14, Question 4], but even without the finiteness conditions, we only know of one construction for finitely generated groups. Using a construction of Gaboriau [Gab00, Proposition VI.16], one can build finitely generated groups with an uncountable range of first $\ell^{2}$-Betti numbers [ITD25, Proposition 5.1].
We present a new construction of groups taking exotic values of second $\ell^{2}$-Betti numbers, and with several additional desirable properties. Given a set $\pi$ of primes, we say that a group $G$ has cohomological dimension $n$ modulo $\pi$, if $G$ has cohomological dimension $n$ over every ring in which each prime from $\pi$ is invertible. Note that if $n<\infty$, this implies that $G$ only contains $\pi$-torsion. The $\ell^{2}$-Euler characteristic of a group, when defined, is the alternating sum of its $\ell^{2}$-Betti numbers.
Theorem A.
Let $p\neq q$ be primes. There exists an uncountable family of groups $(G_{\alpha})_{\alpha\in A}$ and an integer $\chi\geq 1$ with the following properties.
  1. 1.
    $G_{\alpha}$ is simple, has property $(T)$, and is lacunary hyperbolic.
  2. 2.
    $G_{\alpha}$ has cohomological dimension $16$ modulo $\{p,q\}$.
  3. 3.
    The second $\ell^{2}$-Betti numbers $b^{(2)}_{2}(G_{\alpha})$ are pairwise distinct. Further, $b^{(2)}_{n}(G_{\alpha})<\infty$ for every $n\in\mathbb{N}$ and every $\alpha\in A$.
  4. 4.
    The $\ell^{2}$-Euler characteristic is $\chi^{(2)}(G_{\alpha})=b_{2}^{(2)}(G_{\alpha})+\chi$.
By renouncing simplicity, we can also construct groups where a single prime obstructs the finiteness of the cohomological dimension: see Theorem 6.1. Property $(T)$ distinguishes these groups from the examples obtained from Gaboriau’s construction, which arise as amalgamated products [ITD25, Proposition 5.1]. Lacunary hyperbolicity distinguishes them from the counterexamples to the general Atiyah Conjecture, since these all contain lamplighters [Gra16, Problem 3], while lacunary hyperbolic groups cannot (Proposition 4.11).
As noticed by Ioana–Tucker-Drob, it follows easily from Gaboriau’s construction [Gab00, Proposition VI.16] that there exist uncountably many pairwise non-measure equivalent groups [ITD25, Proposition 5.1]. However, in measurable group theory, property $(T)$ is especially desirable, and last year two constructions of measurably diverse families of property $(T)$ groups have appeared. One by the first author and Sun [FFS25], where these groups are distinguished by the whole sequence of their $\ell^{2}$-Betti numbers - in the same vein as the paper by López Neumann [LN23], which constructs an infinite family of finitely presented simple groups. Then very recently by Ioana–Tucker-Drob [ITD25], where the groups have vanishing $\ell^{2}$-Betti numbers, and arise from the technology of wreath-like products [CIOS23]. In particular, these groups have infinite rational cohomological dimension.
Corollary B.
The groups from A are pairwise non-measure equivalent.
Proof.
Let $G_{\alpha}$ and $G_{\beta}$ be groups from the family. If they are measure equivalent with index $c>0$, then $b^{(2)}_{n}(G_{\alpha})=cb^{(2)}_{n}(G_{\beta})$ for all $n\in\mathbb{N}$ by Gaboriau’s theorem [Gab02], and thus $\chi^{(2)}(G_{\alpha})=c\chi^{(2)}(G_{\beta})$. Hence
$cb^{(2)}_{2}(G_{\beta})+\chi=b^{(2)}_{2}(G_{a})+\chi=\chi^{(2)}(G_{\alpha})=c\chi^{(2)}(G_{\beta})=cb^{(2)}_{2}(G_{\beta})+c\chi.$
Since $\chi$ is a positive integer, this shows that $c=1$, and thus $b^{(2)}_{2}(G_{\alpha})=b^{(2)}_{2}(G_{\beta})$, so $\alpha=\beta$. ∎
A version of the construction leading to A produces torsion-free hyperbolic groups with prescribed integer second $\ell^{2}$-Betti number (Theorem 5.1). This has the following consequence.
Theorem C.
Let $(n_{i})_{i\in\mathbb{N}}$ be a sequence of natural numbers. There exist property $(T)$ torsion-free hyperbolic groups $G,G_{i}$ of cohomological dimension $16$, with quotient maps $G\to G_{i}$ such that $G_{i}\xrightarrow{i\to\infty}G$ in the space of marked groups, and
$b^{(2)}_{2}(G_{i})=n_{i}\text{ for all }i\in\mathbb{N}\text{ but }b^{(2)}_{2}(G)=0.$
Therefore the second $\ell^{2}$-Betti number is far from being semi-continuous in the space of marked groups, even with the strong finiteness properties that are enjoyed by torsion-free hyperbolic groups. This is in contrast with the semi-continuity of the first $\ell^{2}$-Betti number [Pic06].
Our computations of $\ell^{2}$-Betti numbers exploit the work of Petrosyan–Sun on group-theoretic Dehn filling [PS24a]. More recently [PS24b], they applied this theory to compute $\ell^{2}$-Betti numbers, by working in convenient classes where the Lück approximation and Atiyah conjectures are known to hold: virtually locally indicable and cocompactly cubulated groups. Note that these are incompatible with property $(T)$.
We instead take a different route, and start from a hyperbolic group enjoying the strong rigidity property $(T_{2})$, introduced by Bader and the second author [BS23, BS25]. This group is a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane (Theorem 3.10), which is the source of the number $16$ in our results. Under this condition, building on an observation from [FF25], we will see that the excision theorem from [PS24a] is sufficient to give exact computations of the second $\ell^{2}$-Betti number of iterated group-theoretic Dehn fillings.
All of the constructions mentioned above, both in this paper and in others, are non-elementary to various extents. Because of this, we feel it is worth presenting possibly the most elementary construction of finitely generated groups achieving uncountably many values of $\ell^{2}$-Betti numbers.
Theorem D.
Let $p$ be a prime. There exists a family of finitely generated groups $(G_{x})_{x\in\mathbb{R}_{>0}}$ such that $b^{(2)}_{1}(G_{x})=0,b^{(2)}_{2}(G_{x})=\infty$, and $b^{(2)}_{3}(G_{x})=x$. Moreover, $G_{x}$ is residually finite and has cohomological dimension $6$ modulo $p$.
Taking a free product with $\mathbb{Z}$ gives an uncountable family of pairwise non-measure equivalent groups (Corollary 7.5). Of course, none of these groups have property $(T)$.
Acknowledgements.
FFF is supported by the Herchel Smith Postdoctoral Fellowship Fund. The authors would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Operators, Graphs, Groups, where work on this paper was undertaken. This work was supported by EPSRC grant EP/Z000580/1.
2 Homology and cohomology
We start by collecting some preliminary facts about homology and cohomology of groups. We fix a ring $R$ for the rest of this section. All rings in this paper are unital.
First is an easy application of the transfer, whose cohomological version is given in [Bro94, Proposition III.10.4 on p. 85]. The homological version can be obtained similarly via the transfer, alternatively it is a direct consequence of the Lyndon–Hochschild–Serre spectral sequence [Bro94, VII.6 on p. 171].
Proposition 2.1.
Let $N<G$ a normal subgroup of finite index, and suppose that $[G:N]$ is invertible in $R$. Then for every $R[G]$-module $M$ and every $n\in\mathbb{N}$, the corestriction induces an isomorphism
$H_{n}(N;M)_{G/N}\cong H_{n}(G;M).$
Next, we look at homology and cohomology of colimit groups, which will be especially relevant for the groups from A.
Proposition 2.2 ([Bie81, Proposition 4.8 on p. 61]).
Let $G_{1}\to G_{2}\to\cdots$ be a directed sequence of groups with colimit $G_{\infty}$, and let $M$ be an $R[G_{\infty}]$-module. Then there is an isomorphism
$H_{n}(G_{\infty};M)\cong\operatorname{colim}H_{n}(G_{i};M).$
This is the standard fact that homology commutes with filtered colimits. This is of course not true in cohomology, but it does work under some additional assumptions. We point out that the following is only used to compute cohomological dimensions of the groups from A and D. If the reader is happy with the standard estimate of cohomological dimension as one more than the homological dimension [Bie81, Theorem 4.6 on p. 60], then they can use Proposition 2.2, and move to the next section.
Proposition 2.3.
Let $G_{1}\to G_{2}\to\cdots$ be a directed sequence of groups with colimit $G_{\infty}$, and let $M$ be an $R[G_{\infty}]$-module. Suppose that the pullback $H^{n}(G_{i+1};M)\to H^{n}(G_{i};M)$ is surjective for all $i\in\mathbb{N}$. Then
[…]
Proof.
A tower of abelian groups
$\cdots\to D_{2}\to D_{1}$
is said to satisfy the Mittag–Leffler condition if for each $k$ there exists $j\geqslant k$ such that for all $i\geqslant j$ the image of $D_{i}\to D_{k}$ equals the image of $D_{j}\to D_{k}$ [Wei94, Definition 3.5.6 on p. 82]. We will replace this with the stronger and easier to check condition that each map $D_{i+1}\to D_{i}$ is surjective.
Let $X_{1}$ be a model of the classifying space $BG_{1}$111also denoted by $K(G_{1},1)$.. Using mapping cones, we may inductively define $X_{i+1}$ as a model of $BG_{i+1}$ such that the map $X_{i}\to X_{i+1}$ is an inclusion of CW-complexes. The colimit $X$, in this case a union, of the $X_{i}$ is a model of $BG_{\infty}$. The universal covering $\widetilde{X}$, which is a free $G_{\infty}$-CW complex, is a model of $EG_{\infty}$. Let $\bar{X}_{i}$ be the restriction of the universal covering to $X_{i}$. By covering theory we have $\bar{X}_{i}\cong G_{\infty}\times_{G_{i}}\widetilde{X}_{i}$. The cellular cochain complexes
$D_{i}=\hom_{R[G_{\infty}]}\bigl(C_{*}(\bar{X}_{i}),M\bigr)\cong\hom_{R[G_{i}]}\bigl(C_{*}(\widetilde{X}_{i}),M\bigr)$
compute the cohomology $H^{*}(G_{i};M)$, and fit into a tower of cochain complexes. Since $\widetilde{X}$ is the colimit of the $G_{\infty}$-CW complexes $\bar{X}_{i}$, the projective limit of the $D_{i}$ is the cochain complex $\hom_{R[G_{\infty}]}\bigl(C_{*}(\widetilde{X}_{i}),M\bigr)$, which computes the cohomology of $G_{\infty}$.
By our choice of $X_{i}$, the maps $D_{i+1}\to D_{i}$ are degree-wise surjective because $C_{\ast}(\bar{X}_{i})$ is degree-wise a direct $R[G_{\infty}]$-summand of $C_{\ast}(\widetilde{X}_{i})$, so we can apply [Wei94, Theorem 3.5.8 on p. 83]222the citation is about homology of chain complexes, hence the left-hand side term being one degree below instead of above and obtain a short exact sequence
[…]
Corollary 2.4.
Let $G_{1}\to G_{2}\to\cdots$ be a directed sequence of groups with colimit $G_{\infty}$. Suppose that the following hold for all $i\in\mathbb{N}$.
  • For every $R[G_{\infty}]$-module $M$ the pullback $H^{n}(G_{i+1};M)\to H^{n}(G_{i};M)$ is surjective.
  • $cd_{R}(G_{i})=n$;
  • $H^{n}(G_{1};R)\neq 0$ and the pullback $H^{n-1}(G_{i+1};R)\to H^{n-1}(G_{i};R)$ is surjective.
Then $cd_{R}(G_{\infty})=n$.
Proof.
By the assumption on surjectivity we can apply Proposition 2.3. First, for every $R[G_{\infty}]$-module $M$:
[…]
thus $cd_{R}(G_{\infty})\leqslant n$. Second,
[…]
and so $cd_{R}(G_{\infty})\geqslant n$. ∎
3 $\ell^{2}$-Betti numbers of tracial unitary representations
We refer the reader to [Lüc02] for details and background, where everything is developed for group von Neumann algebras. Lück formulated his dimension theory for arbitrary finite von Neumann algebras in [Lüc98].
A unitary $G$-representation $\pi$ on a Hilbert space $V_{\pi}$ is called tracial if the associated von Neumann algebra $\mathcal{N}(\pi)\coloneqq\pi(G)^{\prime\prime}$, which is the double commutant of $\pi(G)$ in the algebra of bounded operators on $V_{\pi}$, has a cyclic vector $\xi_{\pi}\in V_{\pi}$ such that
$\tau(T)\coloneqq\langle T\xi_{\pi},\xi_{\pi}\rangle_{V_{\pi}}$
defines a finite trace on $\mathcal{N}(\pi)$. The embedding
$\mathcal{N}(\pi)\hookrightarrow V_{\pi}:T\mapsto T\xi_{\pi}$ (1)
is isometric with respect to the norm $\|T\|^{2}=\tau(T^{\ast}T)$ on $\mathcal{N}(\pi)$ and has dense image. The embedding is naturally a map of left $\mathcal{N}(\pi)$-modules. It extends to an isometric isomorphism of the completion $L^{2}(\mathcal{N}(\pi),\tau)\cong V_{\pi}$. Via this isomorphism, $V_{\pi}$ has a natural right $\mathcal{N}(\pi)$-module structure that commutes with the left $\mathcal{N}(\pi)$-module structure. On the dense set of vectors $T\xi_{\pi}$, the von Neumann algebra $\mathcal{N}(\pi)$ acts from the right via $(T\xi_{\pi})\cdot S=TS\xi_{\pi}$.
We regard the cyclic vector, and therefore $\tau$, as part of the data of a tracial unitary representation. Via the Gelfand–Naimark–Segal (GNS) construction, tracial unitary $G$-representations are in one-to-one correspondence with conjugation-invariant positive functions on $G$ [BdlHV08, Theorem C.4.10 on p. 354].
Example 3.1.
The von Neumann algebra associated to the regular representation $\lambda_{G}$ on $\ell^{2}(G)$ with cyclic vector $1_{G}$ is the group von Neumann algebra of $G$, which we denote by $\mathcal{N}(G)$.
If $G\to Q$ is a surjective group homomorphism, then composing it with $\lambda_{Q}$ yields a tracial unitary $G$-representation on $\ell^{2}(Q)$.
Lück introduced a dimension function
$\dim_{\mathcal{N}}\colon\left\{\text{$\mathcal{N}$-modules}\right\}\to\mathbb{R}_{\geq 0}\cup\{\infty\}$
for arbitrary (algebraic) modules over a finite von Neumann algebra $\mathcal{N}$ with finite normalised trace $\tau$. The choice of $\tau$ is understood to be specified, and we usually drop $\tau$ from the notation. Lück’s dimension function is additive for short exact sequences of $\mathcal{N}$-modules and normalised, i.e. $\dim_{\mathcal{N}}(\mathcal{N})=1$. Moreover:
Lemma 3.2 ([Lüc98, Theorem 0.6 (b) and (c)]).
Let $(V_{i})_{i\in\mathbb{N}}$ be a sequence of $\mathcal{N}$-modules. Then
$\dim_{\mathcal{N}}\left(\bigoplus_{i\in\mathbb{N}}V_{i}\right)=\sum\limits_{i\in\mathbb{N}}\dim_{\mathcal{N}}(V_{i}).$
Let $X$ be a topological $G$-space. We do not require a cellular structure on $X$ nor freeness of the action. Let $\pi$ be a tracial unitary $G$-representation. The equivariant homology
$H_{n}^{G}\bigl(X,\mathcal{N}(\pi)\bigr)=H_{n}\bigl(\mathcal{N}(\pi)\otimes_{\mathbb{Z}[G]}C_{\ast}(X)\bigr)$
is naturally a left $\mathcal{N}(\pi)$-module. Here $C_{\ast}(X)$ denotes the singular chain complex of $X$. We have
$H_{n}(G;\mathcal{N}(\pi))=H_{n}^{G}(EG;\mathcal{N}(\pi)).$
Definition 3.3.
Let $\pi$ be a tracial unitary $G$-representation. The $\ell^{2}$-Betti numbers of $\pi$ are defined as
$b^{(2)}_{n}(G;\pi)=\dim_{\mathcal{N}(\pi)}H_{n}\bigl(G;\mathcal{N}(\pi)\bigr).$
When $\pi$ is the regular representation, we simply write $b^{(2)}_{n}(G)$. When $G\to Q$ is a quotient, and $\pi$ is the pullback of the regular representation of $Q$, we write $b^{(2)}_{n}(G\to Q)$.
Definition 3.4.
Let $\pi$ be a tracial unitary $G$-representation such that $\sum_{n\in\mathbb{N}}b^{(2)}_{n}(G;\pi)<\infty$. Then we define the $\ell^{2}$-Euler characteristic of $\pi$ as
$\chi^{(2)}(G;\pi)=\sum_{n\in\mathbb{N}}(-1)^{n}b^{(2)}_{n}(G;\pi).$
When $\pi$ is the regular representation we just write $\chi^{(2)}(G)$.
The next result generalises the Euler–Poincaré formula for the regular representation to arbitrary tracial unitary representations.
Proposition 3.5.
Let $G$ be a group of type $FP(\mathbb{Q})$. Then the Euler characteristic of $G$ and the $\ell^{2}$-Euler characteristic of every tracial unitary $G$-representation $\pi$ coincide:
$\chi(G)=\chi^{(2)}(G;\pi).$
Proof.
Let $P_{\ast}$ be a finite projective $\mathbb{Q}[G]$-resolution of the trivial $\mathbb{Q}[G]$-module $\mathbb{Q}$. Then $C_{\ast}=\mathcal{N}(\pi)\otimes_{\mathbb{Q}[G]}P_{\ast}$ is a finite projective $\mathcal{N}(\pi)$-chain complex. The proof that
$\sum_{n\geq 0}(-1)^{n}\dim_{\mathcal{N}(\pi)}(C_{n})=\sum_{n\geq 0}(-1)^{n}\dim_{\mathcal{N}(\pi)}\bigl(H_{n}(C_{\ast})\bigr)$
is the same as the Euler–Poincaré formula for the classical Euler characteristic and basically follows from the additivity of the dimension. See the standard proof in Hatcher’s book [Hat02, Theorem 2.44 on p. 146]. ∎
The following result is an analogue of [PT11, Theorem 2.2], which treated the cohomological version, and only for the regular representation. According to the discussion after Eq. 1, the Hilbert space $V_{\pi}$ carries a left $\mathcal{N}(\pi)$-module structure and a right $\mathcal{N}(\pi)$-module structure, in particular, also a right $\mathbb{C}[G]$-module structure. So the homology on the right hand side is naturally a left $\mathcal{N}(\pi)$-module.
Proposition 3.6.
Let $\pi$ be a tracial unitary $G$-representation on the Hilbert space $V_{\pi}$. Then we have
$b^{(2)}_{n}(G;\pi)=\dim_{\mathcal{N}(\pi)}H_{n}\bigl(G;V_{\pi}\bigr)$
for every $n\in\mathbb{N}$.
Proof.
The embedding $\mathcal{N}(\pi)\hookrightarrow V_{\pi}$ of from Eq. 1 is rank dense:
One calls a $\mathcal{N}(\pi)$-linear homomorphism $f\colon M\to N$ of $\mathcal{N}(\pi)$-modules rank dense if for every $x\in N$ there is an increasing sequence of projections $p_{i}\in\mathcal{N}(\pi),i\in\mathbb{N}$, converging to $\mathrm{id}$ such that $p_{i}x\in N$. By [KPV15, Lemma A.5] the embedding $\mathcal{N}(\pi)\hookrightarrow V_{\pi}$ is rank dense because it is dense.
The embedding $\mathcal{N}(\pi)\hookrightarrow V_{\pi}$ is also right $\mathcal{N}(\pi)$-linear, in particular, an embedding of right $\mathbb{C}[G]$-modules, and the right $\mathbb{C}[G]$-action commutes with the left $\mathcal{N}(\pi)$-action. Now we consider the quotient $W\coloneqq V_{\pi}/\mathcal{N}(\pi)$; this is an undesirable object from a functional-analytic perspective, however it makes sense to look at $W$ as an (algebraic) $\mathcal{N}(\pi)$-module. By rank density for every $x\in W$ there is an increasing sequence of projections $p_{i}\in\mathcal{N}(\pi)$ converging to $\mathrm{id}$ such that $p_{i}x=0$ for every $i\in\mathbb{N}$. By the local criterion [Sau05, Theorem 2.4] we obtain that $\dim_{\mathcal{N}(\pi)}(W)=0$. Let $P_{\ast}$ be a free $\mathbb{C}[G]$-resolution of $\mathbb{C}$ of finite type. We have a short exact sequence of $\mathcal{N}(\pi)$-chain complexes
$0\to\mathcal{N}(\pi)\otimes_{\mathbb{C}[G]}P_{\ast}\to V_{\pi}\otimes_{\mathbb{C}[G]}P_{\ast}\to W\otimes_{\mathbb{C}[G]}P_{\ast}\to 0.$
By Lemma 3.2, the right hand side is a chain complex of modules of $\dim_{\mathcal{N}(\pi)}$-dimension zero. By additivity and the long exact sequence in homology, the statement follows. ∎
Proposition 3.6 will be especially useful in conjunction with higher property $(T)$.
Definition 3.7 ([BS23]).
We say that a group $G$ has property $T_{n}]$ if for every $i\leqslant n$ and for every unitary representation $V$, it holds $H^{i}(G;V)=0$. We say that $G$ has property $(T_{n})$ if the vanishing holds under the additional assumption that $V^{G}=0$.
Since we will mostly be working with homology, the following will be more useful.
Proposition 3.8 ([BS25]).
Let $G$ be a group with property $T_{n}]$ (respectively, $(T_{n})$) and of type $FP_{n+1}(\mathbb{Q})$. Then for every $i\leqslant n$ and for every unitary representation $V$ (respectively, every unitary representation $V$ with $V^{G}=0$), it holds $H_{i}(G;V)=0$.
Proof.
Combine [BS25, Lemma 29] and [BS25, Lemma 31]. ∎
Corollary 3.9.
Let $G$ be a group with property $T_{n}]$ (respectively, $(T_{n})$) and of type $FP(\mathbb{Q})$, and let $Q$ be a quotient (respectively, an infinite quotient) of $G$. Then for every $i\leqslant n$ it holds $b^{(2)}_{i}(G\to Q)=0$. Hence
$\chi^{(2)}(G)=\sum\limits_{i>n}(-1)^{i}b^{(2)}_{i}(G\to Q).$
[…]
Proof.
Combine Propositions 3.6 and 3.8. The second statement now follows from Proposition 3.5. ∎
Our starting point will be the following example.
Theorem 3.10 ([BS23]).
Let $G$ be a cocompact lattice in the isometry group of the octonionic hyperbolic plane. Then $G$ has properties $(T_{3})$ and $T_{2}]$.
In the proofs of Theorems A and C, we will actually only need that $G$ has property $(T_{2})$: see Remark 5.3.
Proof.
The ambient Lie group has property $(T_{3})$ by [BS23, Theorem B and Appendix A]. We emphasise that this relies only on a small part of [BS23]: It follows from a result of Kumaresan [Kum80, Theorem 3] for non-trivial irreducible unitary representation of the ambient Lie group, and the ultrapower technique in [BS23]. Finally, property $(T_{3})$ passes to cocompact lattices by [BS23, Lemma 3.13]. Moreover, the second Betti number of $G$ vanishes, because it is the fundamental groups of a non-Hermitian locally symmetric space of rank one [Mat62], so $G$ also has $T_{2}]$. ∎
We conclude with a useful fact about subgroups.
Proposition 3.11 ([Lüc02, Theorem 6.29 on p. 253]).
Let $G\to Q$ be a homomorphism, and let $Q\to L$ be an injective homomorphism. Then
$\dim_{\mathcal{N}(Q)}H_{n}(G;\mathcal{N}(Q))=\dim_{\mathcal{N}(L)}H_{n}(G;\mathcal{N}(L)),$
for all $n\in\mathbb{N}$.
4 Group-theoretic Dehn filling
In this section we introduce group-theoretic Dehn filling, and do some preliminary work towards our main construction.
4.1 Cohen–Lyndon triples
[…]
Definition 4.1.
[…]
$\left<\!\left<N\right>\!\right>=\ast_{t\in T}tNt^{-1}.$
This notion goes back to the work of Cohen–Lyndon on one-relator groups [CL63]; in this form it is defined in [Sun20]. The most important consequence is the following algebraic excision theorem.
Theorem 4.2 ([PS24a, Theorem A (ii)], see also [PS24b, Corollary 4.5]).
Let $(G,P,N)$ be a Cohen–Lyndon triple. Then for all $\mathbb{Z}[\bar{G}]$-modules $M$ and all $n\in\mathbb{N}$, the quotient maps $G\to\bar{G}$ and $P\to\bar{P}$ induce isomorphisms
Hn(G,P;M)Hn(G¯,P¯;M) and $H_{n}(G,P;M)\cong H_{n}(\bar{G},\bar{P};M)\quad\text${ and $\quad$ $\cong H^{n}(\bar{G},\bar{P};M)$.
We will only be concerned with an especially well-behaved case.
Corollary 4.3.
Let $(G,P,N)$ be a Cohen–Lyndon triple. Suppose that $P$ is virtually cyclic, $N$ is infinite cyclic, and hence $\bar{P}$ is finite. Let $R$ be a ring where $[P:N]$ is invertible. Then for all $R[\bar{G}]$-modules $M$ and all $n\geqslant 3$, the quotient map $G\to\bar{G}$ induces isomorphisms
Hn(G;M)Hn(G¯;M) and $H_{n}(G;M)\cong H_{n}(\bar{G};M)\quad\text${ and $\quad$ $\cong H^{n}(\bar{G};M)$.
Moreover, there is an exact sequence
$0\to H_{2}(G;M)\to H_{2}(\bar{G};M)\to M_{\bar{P}}\to H_{1}(G;M).$
Proof.
By the assumption on invertibility of $[P:N]$, we have $cd_{R}(\bar{P})=0$ and $cd_{R}(P)=1$. Therefore $H_{n}(\bar{G},\bar{P};M)\cong H_{n}(\bar{G};M)$ for all $n\geqslant 2$, and $H_{n}(G,P;M)\cong H_{n}(G;M)$ for all $n\geqslant 3$. Using this, for all $n\geqslant 3$, the long exact sequences in homology for the pairs $(\bar{G},\bar{P})$ and $(G,P)$, together with Theorem 4.2, give isomorphisms
$H_{n}(\bar{G};M)\cong H_{n}(\bar{G},\bar{P};M)\cong H_{n}(G,P;M)\cong H_{n}(G;M),$
and similarly for cohomology.
For the second statement, the long exact sequence in homology of the pair $(G,P)$ includes:
$H_{2}(P;M)\to H_{2}(G;M)\to H_{2}(G,P;M)\to H_{1}(P;M)\to H_{1}(G;M).$
The first term vanishes because $cd_{R}(P)=1$. Using $cd_{R}(\bar{P})=0$ and Theorem 4.2, the third term becomes $H_{2}(\bar{G};M)$. Finally, the fourth term can be replaced by $H_{1}(N;M)_{\bar{P}}$ by Proposition 2.1. But $N\cong\mathbb{Z}$ acts trivially on $M$, so $H_{1}(N;M)=M$. ∎
4.2 (Relatively) hyperbolic groups
When $G$ is hyperbolic relative to $P$, group-theoretic Dehn fillings are especially well-behaved [Osi07]. The following was proved in [GMS19], and then in the more general context of groups with hyperbolically embedded subgroups [Sun20].
Theorem 4.4 ([GMS19, Sun20]).
Let $G$ be hyperbolic relative to $P$. Then for every sufficiently deep normal subgroup $N<P$, the triple $(G,P,N)$ is Cohen–Lyndon.
Here a property is said to hold for sufficiently deep normal subgroups $N<P$ if there exists a finite subset $F\subset P\setminus\{1\}$ such that the property holds for all normal subgroups $N<P$ such that $N\cap F=\emptyset$. In this context, group-theoretic Dehn fillings retain many geometric properties.
Theorem 4.5 ([Osi07]).
Let $G$ be hyperbolic relative to $\{P_{\lambda}\}_{\lambda\in\Lambda}$. Let $B\subset G$ be a finite subset. Then for every sufficiently deep normal subgroups $N_{\lambda}<P_{\lambda}$, it holds
  1. 1.
[…]
  2. 2.
[…]
  3. 3.
[…]
We will mostly be concerned with a very particular example of a relatively hyperbolic pair. Let $G$ be a hyperbolic group, and let $g\in G$ be an element of infinite order. Then $g$ is contained in a unique maximal virtually cyclic subgroup denoted $E_{G}(g)$.
Theorem 4.6 ([Bow12, Theorem 7.11]).
Let $G$ be a hyperbolic group, and let $g\in G$ be an element of infinite order. Then $G$ is hyperbolic relative to $E_{G}(g)$.
Putting these results together we obtain the following, which will serve as the inductive step in our constructions.
Proposition 4.7.
[…]
  1. 1.
    The triple $(G,P,N)$ is Cohen–Lyndon;
  2. 2.
    The group $\bar{G}$ is non-elementary hyperbolic with no non-trivial finite normal subgroups;
  3. 3.
    The quotient map $G\to\bar{G}$ is injective when restricted to $B$, and induces an embedding $\bar{P}\hookrightarrow\bar{G}$.
Proof.
By Theorem 4.6, $G$ is hyperbolic relative to $E_{G}(g)$. The Cohen–Lyndon property follows from Theorem 4.4; the hyperbolicity follows from Theorem 4.5, since a group that is hyperbolic relative to a finite subgroup is hyperbolic; the injectivity follows again from Theorem 4.5.
It remains to prove that $\bar{G}$ is non-elementary and without finite normal subgroups. To this end, we find an additional element of infinite order $h\in G$ such that $G$ is hyperbolic relative to $\{E_{G}(g),E_{G}(h)\}$ [Osi07, Lemmas 5.5 and 5.6], see also [Osi06a, Corollaries 1.7 and 4.5]. Consider the normal subgroups $N<E_{G}(g)$ and $\{1\}<E_{G}(h)$, so that the quotient $G\to\bar{G}$ fits into the framework of Theorem 4.5. In particular, $\bar{G}$ is hyperbolic relative to the image of $E_{G}(h)$, which maps injectively and is almost malnormal in $\bar{G}$ by [Osi06b, Theorem 1.4]. If $\bar{G}$ had a non-trivial finite normal subgroup, then a sufficiently high power of $h$ would centralise it, contradicting malnormality. Similarly, if $\bar{G}$ were virtually cyclic, then a sufficiently high power of $h$ would generate a normal subgroup of $\bar{G}$, contradicting malnormality again. ∎
The situation when $E_{G}(g)=\langle g\rangle$ is especially convenient. Following [OT13], we call such elements special.
Proposition 4.8 ([OT13, Proposition 3.4]).
[…]
In the cases above, to obtain desireable properties we needed to pass to a sufficiently deep normal subgroup of the elementary closure of an element. Let us highlight one situation in which this is not necessary.
Lemma 4.9.
[…]
  1. 1.
    The triple $(G,\langle g\rangle,\langle g\rangle)$ is Cohen–Lyndon;
  2. 2.
    The group $\bar{G}$ is torsion-free non-elementary hyperbolic;
  3. 3.
    The quotient map $G\to\bar{G}$ is injective when restricted to $B$.
Proof.
This obtained in the course of the proof of [FFS25, Theorem 3.1], the second and third statement are consequences of the theorem, the first one is [FFS25, Lemma 3.5]. ∎
4.3 Lacunary hyperbolic groups
A group is said to be lacunary hyperbolic if one of its asymptotic cones is an $\mathbb{R}$-tree. We will use the following useful characterisation.
Theorem 4.10 ([OOS09, Theorem 1.1]).
Let $G_{1}\to G_{2}\to\cdots$ be a directed sequence of epimorphisms of hyperbolic groups, with colimit $G_{\infty}$. Let $S_{1}$ be a finite generating set for $G_{1}$, and let $S_{i}$ be its pushforward to $G_{i}$; let $\delta_{i}$ be the hyperbolicity constant of $\mathrm{Cay}(G_{i},S_{i})$. For each $i\in\mathbb{N}$, let $\rho_{i}$ be the radius of the largest ball centred at the identity on which the quotient $G_{i}\to G_{i+1}$ is injective.
If $\delta_{i}=\mathrm{o}(\rho_{i})$, then $G_{\infty}$ is lacunary hyperbolic.
As we mentioned in the introduction, many earlier constructions of exotic values of von Neumann dimensions associated to groups involved lamplighters. Hence we record the following:
Proposition 4.11 ([OOS09]).
Let $H$ be a non-trivial group, then $H\wr\mathbb{Z}$ cannot be a subgroup of a lacunary hyperbolic group.
Proof.
Up to passing to a subgroup, it suffices to show this for $H=\mathbb{Z}$ or $H=\mathbb{Z}/p$ for a prime $p$. When $H=\mathbb{Z}$, this follows from the more general fact that a lacunary hyperbolic group cannot contain a copy of $\mathbb{Z}^{2}$ by [OOS09, Corollary 3.21(a)]. For $H=\mathbb{Z}/p$, it follows from [OOS09, Corollary 3.21(c)] and its proof, see [OOS09, Remark 3.22]. ∎
5 Failure of semi-continuity
In this section we prove C. We start with a more precise statement.
Theorem 5.1.
Let $G$ be a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane, and let $B\subset G$ be a finite subset. Then for every $n\in\mathbb{N}$, there exists a torsion-free hyperbolic quotient $G\to G_{n}$, injective on $B$, of cohomological dimension $16$, such that for every unitary $G_{n}$-representation $V$, the second homology is $H_{2}(G_{n};V)\cong V^{n}$.
Cocompact lattices in the isometry group of the octonionic hyperbolic plane exist [BHC62], and they can be chosen torsion-free up to passing to a finite index subgroup [Sel60].
Proof.
The group $G$ is torsion-free hyperbolic of cohomological dimension $16$, and by Theorem 3.10 it has property $T_{2}]$. We construct $G_{n}$ by induction on $n$, starting with $G_{0}=G$, which works by Proposition 3.8.
Suppose by induction that $G_{n-1}$ has been constructed. Lemma 4.9 gives an element $g\in G_{n-1}$ and we define $G_{n}$ to be the corresponding group-theoretic Dehn filling. Lemma 4.9 gives injectivity on $B$ and the fact that $G_{n}$ is torsion-free hyperbolic. The first part of Corollary 4.3 implies that the cohomological dimension of $G_{n}$ is $16$. Moreover, since $G_{n-1}$ has property $(T)$, applying Proposition 3.8 and the second part of Corollary 4.3, we obtain, for every unitary $G_{n}$-representation $V$, a short exact sequence
$0\to H_{2}(G_{n-1};V)\to H_{2}(G_{n};V)\to V\to 0.$
By the induction hypothesis the first term is $V^{n-1}$. Therefore $H_{2}(G_{n};V)\cong V^{n}$, which proves the claim and hence the theorem. ∎
Remark 5.2.
The groups $G_{n}$ of course fail to have property $(T_{2})$, despite being quotients of the $(T_{2})$ group $G$, as remarked in [FF25]. Still, the second homology (and similarly the second cohomology) functor enjoys some strong form of rigidity: in a way, its output can be read directly from its input. It would be interesting to explore whether a (co)homological rigidity of this form characterises quotients of $(T_{2})$ groups.
Proof of C.
We have $b^{(2)}_{2}(G)=0$ by Corollary 3.9. For each $i\in\mathbb{N}$, we let $B_{i}$ be the ball of radius $i$ in $G$, for a fixed choice of finite generating set, and $n_{i}\in\mathbb{N}$ as in the statement of C. Then Theorem 5.1 builds a group $G_{i}$, which by Proposition 3.6 has
$b^{(2)}_{2}(G_{i})=\dim_{\mathcal{N}(G_{i})}H_{2}(G_{i};\ell^{2}(G_{i}))=\dim_{\mathcal{N}(G_{i})}\ell^{2}(G_{i})^{n_{i}}=n_{i}.$
Since $G\to G_{i}$ is injective on larger and larger balls, $G_{i}\xrightarrow{i\to\infty}G$ in the space of marked groups. The groups $G_{i}$ have property $(T)$ being quotients of $G$. The other properties of $G_{i}$ are all contained in the statement of Theorem 5.1. ∎
Remark 5.3.
In Theorem 5.1, we used the full power of property $T_{2}]$, but the proof shows that if $V$ is assumed to have no non-zero invariant vectors, then property $(T_{2})$ suffices. In the proof of C, we applied Theorem 5.1 only with $V=\ell^{2}(G_{i})$, which has no non-zero invariant vectors, so property $(T_{2})$ would have been sufficient. Similarly, the computation of second $\ell^{2}$-Betti numbers in the next section could be obtained only using property $(T_{2})$.
6 Proof of the main theorem
In this section we prove A. Because the proof has many moving parts, we start with a simpler version, that only involves one prime, and does not give simplicity.
Theorem 6.1.
Let $p$ be a prime. There exists an uncountable family of groups $(G_{\alpha})_{\alpha\in A}$ and an integer $\chi\geq 1$ with the following properties.
  1. 1.
    $G_{\alpha}$ has property $(T)$ and is lacunary hyperbolic.
  2. 2.
    $G_{\alpha}$ has cohomological dimension $16$ modulo $p$.
  3. 3.
    The second $\ell^{2}$-Betti numbers $b^{(2)}_{2}(G_{\alpha})$ are pairwise distinct.
  4. 4.
    The $\ell^{2}$-Euler characteristic is $\chi^{(2)}(G_{\alpha})=b_{2}^{(2)}(G_{\alpha})+\chi$.
Proof.
We choose as an index set $A\coloneqq\{2,3\}^{\mathbb{N}}$, and fix an element $\alpha\in A$. The group $G_{\alpha}$ will arise as the colimit of a direct sequence of epimorphisms of hyperbolic groups $G_{1}\to G_{2}\to\cdots$, where $G_{1}$ is a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane, which has Euler characteristic $\chi\geqslant 1$ [Lüc02, Corollary 5.16(1) on p. 231]. We fix a finite generating set $S_{1}$ for $G_{1}$ and let $S_{i}$ be the pushforward to $G_{i}$. We also specify elements $g_{i}\in G_{i}$, and will prove the following additional properties.
  1. 1.
    $G_{i}$ is non-elementary hyperbolic with no non-trivial finite normal subgroups.
  2. 2.
    $\mathrm{Cay}(G_{i},S_{i})$ is $\delta_{i}$-hyperbolic, and the quotient $G_{i}\to G_{i+1}$ is injective on the ball of radius $i\delta_{i}$ centred at the identity.
  3. 3.
    Each quotient $G_{i}\to G_{i+1}$ induces isomorphisms in homology and cohomology with coefficients in $R[G_{i+1}]$-modules and degree at least $3$, where $R$ is any ring where $p$ is invertible.
  4. 4.
    The element $g_{i}\in G_{i}$ has infinite order, and for all $j>i$ its image in $G_{j}$ has order $p^{a_{i}}$, where $a_{i}\in\mathbb{N}$ is a sequence with the property that $a_{i+1}-a_{i}-1$ is a positive power of $\alpha(i)\in\{2,3\}$.
  5. 5.
    For every unitary $G_{i}$-representation $V$, the second homology is
    $H_{2}(G_{i};V)=\bigoplus_{j<i}V_{g_{j}}.$
Let us first prove that the colimit $G_{\alpha}$ has the desired properties. The group $G_{1}$ has property $(T)$ by Theorem 3.10, hence so does $G_{\alpha}$, being its quotient. Item 2 implies that $G_{\alpha}$ is lacunary hyperbolic, by Theorem 4.10. Now $G_{1}$ is the fundamental group of a closed $16$-dimensional aspherical manifold, so it has cohomological dimension $16$, and $H^{16}(G_{1};R)\cong R$ for every ring $R$. Therefore Item 3 and Corollary 2.4 imply that $G_{\alpha}$ has cohomological dimension $16$ modulo $p$.
Next, we compute $b^{(2)}_{2}(G_{\alpha})$. By Proposition 2.2, we have
$H_{2}(G_{\alpha};\ell^{2}(G_{\alpha}))\cong\bigoplus_{i\geqslant 1}\ell^{2}(G_{\alpha})_{g_{i}}.$
Each $g_{i}$ has order $p^{a_{i}}$ in $G_{\alpha}$, hence by Proposition 3.6 and Proposition 3.11:
$\dim_{\mathcal{N}(G_{\alpha})}\ell^{2}(G_{\alpha})_{g_{i}}$ $=\dim_{\mathcal{N}(G_{\alpha})}H_{0}(\langle g_{i}\rangle,\ell^{2}(G_{\alpha}))=\dim_{\mathcal{N}(G_{\alpha})}H_{0}(\langle g_{i}\rangle,\mathcal{N}(G_{\alpha}))$
$=\dim_{\mathcal{N}(\langle g_{i}\rangle)}H_{0}(\langle g_{i}\rangle,\mathcal{N}(\langle g_{i}\rangle))=b^{(2)}_{0}(\langle g_{i}\rangle)=p^{-a_{i}}.$
By Proposition 3.6 and Lemma 3.2, we therefore have
$b^{(2)}_{2}(G_{\alpha})=\dim_{\mathcal{N}(G_{\alpha})}H_{2}(G_{\alpha};\ell^{2}(G_{\alpha}))=\sum\limits_{i\geqslant 1}\frac{1}{p^{a_{i}}}.$
Reading this number in base $p$, we see a $1$ at position $a_{i}$, then a string of 0s of length $(a_{i+1}-a_{i}-1)$ before the next $1$ at position $a_{i+1}$. The length of this string of 0s is either a power of $2$ or a power of $3$ depending on whether $\alpha(i)=2$ or $3$. This shows that we can read $\alpha$ from $b^{(2)}_{2}(G_{\alpha})$.
Finally, we compute $\chi^{(2)}(G_{\alpha})$. By Item 3 and Proposition 2.2, for all $n\geqslant 3$ we have
$b^{(2)}_{n}(G_{\alpha})=\dim_{\mathcal{N}(G_{\alpha})}H_{n}(G_{\alpha};\mathcal{N}(G_{\alpha}))=\dim_{\mathcal{N}(G_{\alpha})}H_{n}(G_{1};\mathcal{N}(G_{\alpha}))=b^{(2)}_{n}(G_{1}\to G_{\alpha}).$
Since $G_{\alpha}$ has property $(T)$, $b^{(2)}_{1}(G_{\alpha})=0$ by [BV97, Corollary 6]. Hence
$\chi^{(2)}(G_{\alpha})$ $=\sum\limits_{n\in\mathbb{N}}(-1)^{n}b^{(2)}_{n}(G_{\alpha})=b^{(2)}_{2}(G_{\alpha})+\sum\limits_{n\geqslant 3}(-1)^{n}b^{(2)}_{n}(G_{\alpha})$
$=b^{(2)}_{2}(G_{\alpha})+\sum\limits_{n\geqslant 3}(-1)^{n}b^{(2)}_{n}(G_{1}\to G_{\alpha})=b^{(2)}_{2}(G_{\alpha})+\chi,$
where the last equality follows from Corollary 3.9.
[…]
By Proposition 4.7 again, the triple $(G_{i},\langle g_{i}\rangle,\langle g_{i}^{k}\rangle)$ is Cohen–Lyndon. Because $k$ is a power of $p$, for every ring $R$ in which $p$ is invertible, we can apply Corollary 4.3. The first part immediately gives Item 3. Moreover, the second part, together with the fact that $G_{i}$ has property $(T)$, gives a short exact sequence
$0\to H_{2}(G_{i};V)\to H_{2}(G_{i+1};V)\to V_{g_{i}}\to 0.$
The induction hypothesis gives
$H_{2}(G_{i+1};V)=\bigoplus\limits_{j\leqslant i}V_{g_{j}},$
proving Item 5. This concludes the construction, hence the proof. ∎
We now prove A. The proof will be similar, but more technical, in that we need to leverage an additional prime, and the full power of Proposition 4.8, to make the group $G_{\alpha}$ simple. This improvement employs a trick from [OT13], where it was used to construct simple groups with large first $\ell^{2}$-Betti numbers. The following will be our starting point.
Lemma 6.2.
Let $G$ be a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane. Let $q$ be a prime. Then there exists a quotient $G\to G_{1}$ with the following properties.
  1. 1.
    $G_{1}$ is non-elementary hyperbolic with no non-trivial finite normal subgroups.
  2. 2.
    The quotient $G\to G_{1}$ induces isomorphisms in homology and cohomology with coefficients in $R[G_{1}]$-modules and degree at least $3$, where $R$ is any ring where $q$ is invertible.
  3. 3.
    $G_{1}$ is generated by elements $\{x_{1},\ldots,x_{m}\}$, each of which has order a power of $q$.
  4. 4.
    There is a finite set $\mathcal{E}$ of finite subgroups of $G_{1}$, such that for every unitary $G_{1}$-representation $V$, the second homology is
    $H_{2}(G_{1};V)=\bigoplus_{E\in\mathcal{E}}V_{E}.$
Proof.
We start by choosing a generating set $\{x_{1}^{0},\ldots,x_{m}^{0}\}$ of $G$. For $i=0,\ldots,m$, we construct by induction a quotient $G\to G_{i/m}$ with the following properties.
  1. 1.
    $G_{i/m}$ is non-elementary hyperbolic with no non-trivial finite normal subgroups.
  2. 2.
    The quotient $G\to G_{i/m}$ induces isomorphisms in homology and cohomology with coefficients in $R[G_{i/m}]$-modules and degree at least $3$, where $R$ is any ring where $q$ is invertible. In particular $G_{i/m}$ has cohomological dimension $16$ modulo $q$, and so it only contains $q$-torsion [Bie81, Proposition 4.11 on p. 63].
  3. 3.
    $G_{i/m}$ is generated by elements $\{x_{1}^{i},\ldots,x_{m}^{i}\}$, where $x_{1}^{i},\ldots,x_{i}^{i}$ have order a power of $q$.
  4. 4.
    There is a finite set of finite subgroups $\mathcal{E}_{i}$ of $G_{i/m}$, such that for every unitary $G_{i/m}$-representation $V$, the second homology is
    $H_{2}(G_{i/m};V)=\bigoplus_{E\in\mathcal{E}_{i}}V_{E}.$
Then the final group $G_{m/m}$ will be the group $G_{1}$ in the statement of the lemma, with $\mathcal{E}=\mathcal{E}_{m}$ and $x_{j}=x_{j}^{m}$. The base case for $i=0$ is $G$ itself, with $\mathcal{E}=\emptyset$, where we do not need to check anything for the first three items, and Item 4 follows from Proposition 3.8 and Theorem 3.10.
Suppose by induction that $G_{i/m}$ has been constructed. If $x_{i+1}^{i}$ is torsion, then it is $q$-torsion, so we set $G_{(i+1)/m}=G_{i/m}$, $x^{i+1}_{j}=x^{i}_{j}$ for all $j$, and $\mathcal{E}_{i+1}=\mathcal{E}_{i}$. Otherwise, consider the infinite virtually cyclic group $E_{G_{i/m}}(x_{i+1})$. This has a unique maximal finite normal subgroup of order a power of $q$, and the quotient is isomorphic to $\mathbb{Z}$ (or possibly to $D_{\infty}$, in case $q=2$). We replace $x_{i+1}^{i}$ with a root $x^{i+1}_{i+1}$ that generates the $\mathbb{Z}$ quotient (or the index-$2$ subgroup of the $D_{\infty}$ quotient), so that $E_{G_{i/m}}(x^{i+1}_{i+1}):\langle x^{i+1}_{i+1}\rangle]$ is a power of $q$. We now let $G_{(i+1)/m}$ be the group-theoretic Dehn filling for the triple $(G_{i/m},E_{G_{i/m}}(x^{i+1}_{i+1}),\langle(x^{i+1}_{i+1})^{k}\rangle)$, where $k$ is a sufficiently large power of $q$. By Proposition 4.7, we see that $G_{(i+1)/m}$ satisfies Items 1 and 3, choosing $x^{i+1}_{j}=x^{i}_{j}$ for all $j\neq i+1$. We also choose $k$ to be large enough so that the orders of $x_{1}^{i},\ldots,x_{i}^{i}$ and their images $x_{1}^{i+1},\ldots,x_{i}^{i+1}$ have the same order, and each finite subgroup from $\mathcal{E}_{i}$ embeds into $G_{(i+1)/m}$. Moreover, by Proposition 4.7 again, this triple is Cohen–Lyndon, so Corollary 4.3 applies, and proves Item 2. Finally, let $E\coloneqq E_{G_{i/m}}(x_{i+1}^{i+1})/\langle(x_{i+1}^{i+1})^{k}\rangle$. Corollary 4.3 and property $(T)$ for $G_{i/m}$ give a short exact sequence
$0\to H_{2}(G_{i/m};V)\to H_{2}(G_{(i+1)/m};V)\to V_{E}\to 0,$
so setting $\mathcal{E}_{i+1}=\mathcal{E}_{i}\cup\{E\}$, we obtain Item 4 and conclude. ∎
Proof of A.
We proceed as in the proof of Theorem 6.1, using the prime $p\neq q$, but starting with the group $G_{1}$ from Lemma 6.2. The result will be a group $G_{\alpha}$ with the same properties as in Theorem 6.1, with the exception that cohomological dimension now has to be taken modulo $\{p,q\}$, and $b^{(2)}_{2}(G_{\alpha})$ is computed starting from
$b^{(2)}_{2}(G_{1}\to G_{\alpha})=\dim_{\mathcal{N}(G_{\alpha})}\Bigl(\bigoplus_{E\in\mathcal{E}}\mathcal{N}(G_{\alpha})_{E}\Bigr)=\sum\limits_{E\in\mathcal{E}}|E|^{-1}.$
Since this starting value is independent of $\alpha$, the same computations go through. For the finiteness of the $\ell^{2}$-Betti numbers of $G_{\alpha}$ we refer to Remark 6.3.
In each induction step in the proof of Theorem 6.1, we chose arbitrary special elements $g_{i}\in G_{i}$. We will now explain how to use the full power of Proposition 4.8 in the construction so that this time the colimit is simple. Enumerate the set $\mathcal{X}\coloneqq\{x_{1},\ldots,x_{m}\}\times G_{1}$. When the group $G_{i}$ has been built, let $(x,h)$ denote the $i$-th element of $\mathcal{X}$. If $h$ is trivial, then we pick $g_{i}$ to be an arbitrary special element, and proceed as before. Otherwise, we use Proposition 4.8 to pick $g_{i}\in x^{-1}\left<\!\left<h\right>\!\right>_{G_{i}}$. We claim that now $x\in\left<\!\left<h\right>\!\right>_{G_{i+1}}\leqslant G_{i+1}$. Indeed, consider the image of $x$ in the quotient $G_{i+1}/\left<\!\left<h\right>\!\right>_{G_{i+1}}$. On the one hand, it has order a power of $q$, because $x$ has order a power of $q$ in $G_{1}$. On the other hand, it equals the image of $g_{i}$, which has order a power of $p$ in $G_{i+1}$, by construction of the quotient $G_{i}\to G_{i+1}$. Because $p\neq q$, we see that $x$ must be trivial in $G_{i+1}/\left<\!\left<h\right>\!\right>_{G_{i+1}}$, which proves our claim.
As a result, for every $(x,h)\in\mathcal{X}$, either $h$ is trivial in $G_{\alpha}$, or $x$ lies in the normal closure of $h$ in $G_{\alpha}$. Because $x$ ranges over a generating set for $G_{1}$ and $h$ ranges over $G_{1}$, this shows that every non-trivial element of $G_{\alpha}$ normally generates, that is, $G_{\alpha}$ is simple. ∎
Remark 6.3.
Note that the constructions give some information about higher $\ell^{2}$-Betti numbers as well. Indeed, $G_{\alpha}$ is obtained from a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane, denoted $G$, by a sequence of group-theoretic Dehn fillings satisfying Corollary 4.3. Using this and Proposition 2.2 we have
$b^{(2)}_{n}(G_{\alpha})=b^{(2)}_{n}(G\to G_{\alpha})$
for all $n\geqslant 3$. Since $G$ is of type $F$, this implies that all $\ell^{2}$-Betti numbers of $G_{\alpha}$ are finite. We know from Theorem 3.10 that $G$ has property $(T_{3})$, so Corollary 3.9 gives $b^{(2)}_{3}(G_{\alpha})=0$.
Moreover, $G$ satisfies the Singer conjecture [Lüc02, Corollary 5.16 on p. 231], that is $b^{(2)}_{n}(G)=0$ for all $n\neq 8$ and $b^{(2)}_{8}(G)=\chi\geqslant 1$. This does not give information about $b^{(2)}_{n}(G\to G_{\alpha})$, however, assuming Lück’s approximation conjecture [Lüc02, Chapter 13], and repeating the construction ensuring that $G\to G_{\alpha}$ is injective on a large ball centred at the identity, we would have that $b^{(2)}_{n}(G\to G_{\alpha})$ is close to $b^{(2)}_{n}(G)$. It would then follow that $b^{(2)}_{n}(G_{\alpha})$ is close to 0, for $n\geqslant 4,n\neq 8$, and close to $\chi$ for $n=8$.
Remark 6.4.
Taking $V=\mathbb{C}$ to be the trivial unitary representation, the proofs of Theorems 6.1 and A, and Proposition 2.2, show that $H_{2}(G_{\alpha};\mathbb{C})$ is infinite-dimensional. It follows that the groups $G_{\alpha}$ do not have type $FP_{2}(\mathbb{Q})$.
7 An elementary construction
We conclude with the elementary construction giving D. We start with two basic product formulas.
Proposition 7.1 ([Lüc02, Theorem 6.54(5) on p. 266]).
Let $G_{1},G_{2}$ be two groups. Then
$b^{(2)}_{n}(G_{1}\times G_{2})=\sum\limits_{i+j=n}b^{(2)}_{i}(G_{1})b^{(2)}_{j}(G_{2}).$
Proposition 7.2.
Let $(G_{i})_{i\in\mathbb{N}}$ be infinite groups, and let $G$ be their free product. Then $b^{(2)}_{1}(G)=\infty$ and
$b^{(2)}_{n}(G)=\sum\limits_{i\in\mathbb{N}}b^{(2)}_{n}(G_{i})$
for all $n\geqslant 2$.
This is usually stated for finite free products only [Lüc02, Theorem 1.35(5) on p. 38] so we include a proof.
Proof.
Let $P_{i}$ denote the free product $G_{1}\ast\cdots\ast G_{i}$, so $P_{i+1}=P_{i}\ast G_{i+1}$. By Mayer–Vietoris, for all $n\geqslant 2$ we have an isomorphism
$H_{n}(P_{i+1};\mathcal{N}(G))\cong H_{n}(P_{i};\mathcal{N}(G))\oplus H_{n}(G_{i+1};\mathcal{N}(G)),$
which shows by induction that
$H_{n}(P_{i};\mathcal{N}(G))=\bigoplus_{j\leqslant i}H_{n}(G_{j};\mathcal{N}(G)).$
By Mayer–Vietoris again, we have an exact sequence
$0\to H_{1}(P_{i+1};\mathcal{N}(G))\to H_{1}(P_{i};\mathcal{N}(G))\oplus H_{1}(G_{i+1};\mathcal{N}(G))\to\mathcal{N}(G)\to\mathcal{N}(G)_{P_{i+1}}.$
By Proposition 3.11, the last term has dimension
$\dim_{\mathcal{N}(G)}H_{0}(P_{i+1};\mathcal{N}(G))=\dim_{\mathcal{N}(P_{i+1})}H_{0}(P_{i+1};\mathcal{N}(P_{i+1}))=b^{(2)}_{0}(P_{i+1})=0.$
So by induction
$H_{1}(P_{i};\mathcal{N}(G))=\left(\bigoplus_{j\leqslant i}H_{1}(G_{j};\mathcal{N}(G))\right)\oplus V_{i},$
where $V_{i}$ is a submodule of $\mathcal{N}(G)^{i-1}$ of codimension 0. We conclude from Propositions 2.2 and 3.11, together with the fact that $\dim_{\mathcal{N}(G)}$ respects directed unions [Lüc98, Theorem 0.6(b)]. ∎
We can now begin the construction.
Proposition 7.3.
Let $p$ be a prime. There exists a family of groups $(C_{x})_{x\in\mathbb{R}_{>0}}$ with the following properties.
  1. 1.
    Each $C_{x}$ is countable and residually finite.
  2. 2.
    $b^{(2)}_{1}(C_{x})=\infty,b^{(2)}_{2}(C_{x})=x$.
  3. 3.
    Each $C_{x}$ has cohomological dimension $2$ modulo $p$. In particular, $b^{(2)}_{n}(C_{x})=0$ for all $n\geqslant 3$.
Proof.
Write $x$ in base-$p$ as
$x=\sum_{i\in\mathbb{N}}\frac{r_{i}}{p^{i}},$
where $r_{0}\in\mathbb{N}$; $r_{i}\in\{0,1,\ldots,p-1\}$ for $i\geqslant 1$. For elements in $\mathbb{Z}\left[\frac{1}{p}\right]$ we pick the periodic representation, which allows to have infinite sums for every $x$. Now for each $i$ such that $r_{i}\neq 0$ we let $A_{i}\coloneqq\mathbb{Z}/p^{i}\mathbb{Z}\times F_{r_{i}+1}^{2}$, where $F_{r}$ is the free group of rank $r$. Then
$b^{(2)}_{0}(A_{i})=b^{(2)}_{1}(A_{i})=0;b^{(2)}_{2}(A_{i})=\frac{r_{i}}{p^{i}},$
by Proposition 7.1. Moreover, $A_{i}$ has cohomological dimension $2$ modulo $p$.
Now let
$C_{x}\coloneqq\Asterisk_{i\geqslant 0}A_{i}.$
By Proposition 7.2, we have the desired values of $\ell^{2}$-Betti numbers. Moreover, $C_{x}$ has cohomological dimension $2$ modulo $p$, by Corollary 2.4. Finally, $C_{x}$ is residually finite, being a free product of residually finite groups. ∎
Proof of D.
Let $C_{x}$ be the groups from Proposition 7.3. We embed each $C_{x}$ into a finitely generated residually finite group $R_{x}$ following [WZ96]. The group $R_{x}$ is an HNN extension of $F_{2}\times C_{x}$ over a free subgroup, therefore it has cohomological dimension $3$ modulo $p$. Moreover, $C_{x}$ is separable, that is closed in the profinite topology [WZ96, Lemma 4].
Consider the group $R_{x}\times F_{2}^{3}$: it has cohomological dimension $6$ modulo $p$, and contains $C_{x}$ as a separable subgroup. Moreover, $b^{(2)}_{n}(R_{x}\times F_{2}^{3})=0$ for $n\leqslant 3$, by Proposition 7.1. Define
$G_{x}\coloneqq(R_{x}\times F_{2}^{3})\Asterisk_{C_{x}}(R_{x}\times F_{2}^{3}),$
where the double is over the copy of $C_{x}$ coming from the first coordinate of each factor. Because $C_{x}$ is separable, $G_{x}$ is residually finite, and moreover it has cohomological dimension $6$ modulo $p$. By Mayer–Vietoris:
$H_{n}((R_{x}\times F_{2}^{3});\mathcal{N}(G_{x}))^{2}\to H_{n}(G_{x};\mathcal{N}(G_{x}))\to H_{n-1}(C_{x};\mathcal{N}(G_{x}))\to H_{n-1}(R_{x}\times F_{2}^{3});\mathcal{N}(G_{x}))^{2}.$
Taking $\dim_{G_{x}}$ and $n\leqslant 3$, the first and last term give $2b^{(2)}_{n}(R_{x}\times F_{2}^{3})=2b^{(2)}_{n-1}(R_{x}\times F_{2}^{3})=0$, by Proposition 3.11, so
$b^{(2)}_{n}(G_{x})=b^{(2)}_{n-1}(C_{x}).$
The value of $\ell^{2}$-Betti numbers are as desired, by Proposition 7.3. ∎
Remark 7.4.
The construction is quite flexible, and can be made more elementary, or less elementary, depending on what properties one wants to impose, if we forgo residual finiteness.
  1. 1.
    We can choose $R_{x}$ to be a finitely generated group that is an HNN extension of $F_{2}\ast C_{x}$ over a free subgroup, therefore has cohomological dimension $2$ modulo $p$ [HNN49]. This is the first and easiest embedding result for countable groups into finitely generated groups. Then $R_{x}\times F_{2}^{3}$ has cohomological dimension $5$ modulo $p$, hence so does $G_{x}$.
  2. 2.
    We can bring the dimension down by one more, up to possibly adding some primes in the torsion. Let $L$ be a generalised triangle group that is hyperbolic and has property $(T)$ [CCKW22]. Then there exists a finite set of primes $\pi$ (the ones appearing in the orders of the finite vertex groups) such that $L$ has cohomological dimension $2$ modulo $\pi$. Using [FFS25, Theorem 4.1], we can embed $C_{x}$ into a group $R_{x}$ that is a quotient of $L$ and has cohomological dimension $2$ modulo $\pi\cup\{p\}$. In particular $R_{x}$ has property $(T)$ and so $b^{(2)}_{1}(R_{x})=0$ by [BV97, Corollary 6]. So in the construction above it suffices to take $R_{x}\times F_{2}^{2}$ to get the correct vanishing of $\ell^{2}$-Betti numbers, and thus the resulting $G_{x}$ has cohomological dimension $4$ modulo $\pi\cup\{p\}$.
  3. 3.
    If we forgo the finiteness of the dimension, then we can gain control on all other $\ell^{2}$-Betti numbers. We first choose $R_{x}^{\prime}$ to be any finitely generated group containing $C_{x}$, and then let $R_{x}\coloneqq R_{x}^{\prime}\wr\mathbb{Z}$, which has vanishing $\ell^{2}$-Betti numbers in all degrees [WWZZ25, Remark 7.6]. So in the construction above we do not need to take products with $F_{2}$, and directly get $b^{(2)}_{n}(G_{x})=b^{(2)}_{n-1}(C_{x})$ for all $n\in\mathbb{N}$. Therefore $b^{(2)}_{n}(G_{x})=0$ for all $n\geqslant 4$.
Corollary 7.5.
Let $(G_{x})_{x\in\mathbb{R}_{>0}}$ be as in D. Then the groups $G_{x}\ast\mathbb{Z}$ are pairwise non-measure equivalent.
Proof.
If $G_{x}\ast\mathbb{Z}$ and $G_{y}\ast\mathbb{Z}$ are measure equivalent with index $c>0$, then $b^{(2)}_{n}(G_{x}\ast\mathbb{Z})=cb^{(2)}_{n}(G_{y}\ast\mathbb{Z})$ for all $n\in\mathbb{N}$ by Gaboriau’s theorem [Gab02]. So $1=b^{(2)}_{1}(G_{x}\ast\mathbb{Z})=cb^{(2)}_{1}(G_{y}\ast\mathbb{Z})=c$, that is $c=1$, and so $x=b^{(2)}_{3}(G_{x}\ast\mathbb{Z})=cb^{(2)}_{3}(G_{y}\ast\mathbb{Z})=cy=y$. ∎

### v2

@@
-Kazhdan groups of dimension $16$ with an uncountable range of second $\ell^{2}$-Betti numbers
+Kazhdan groups of dimension $16$ with prescribed second $\ell^{2}$-Betti number
 Abstract
-We construct a family of simple, lacunary hyperbolic groups with property $(T)$ that have rational cohomological dimension $16$ and uncountably many values of second $\ell^{2}$-Betti numbers.
+We construct a family of simple, lacunary hyperbolic groups with property $(T)$ that have rational cohomological dimension $16$ and whose second $\ell^{2}$-Betti number can be prescribed to be any positive real.
+Moreover, we construct hyperbolic groups with property $(T)$ whose second $\ell^{2}$-Betti number can be prescribed to be any non-negative rational.
 Along the way, we present new constructions of measurably diverse finitely generated groups, and we prove that the second $\ell^{2}$-Betti number is far from being semi-continuous in the space of marked groups, even assuming good finiteness properties.
@@
 Theorem A.
-Let $p\neq q$ be primes. There exists an uncountable family of groups $(G_{\alpha})_{\alpha\in A}$ and an integer $\chi\geq 1$ with the following properties.
+There exists an integer $\chi\geqslant 1$ such that the following holds. Let $p\neq q$ be primes. There exists a family of groups $(G_{x})_{x\in\mathbb{R}_{>0}}$ with the following properties.
   1. 1.
-    $G_{\alpha}$ is simple, has property $(T)$, and is lacunary hyperbolic.
+    $G_{x}$ is simple, has property $(T)$, and is lacunary hyperbolic.
   2. 2.
-    $G_{\alpha}$ has cohomological dimension $16$ modulo $\{p,q\}$.
+    $G_{x}$ has cohomological dimension $16$ modulo $\{p,q\}$.
   3. 3.
-    The second $\ell^{2}$-Betti numbers $b^{(2)}_{2}(G_{\alpha})$ are pairwise distinct. Further, $b^{(2)}_{n}(G_{\alpha})<\infty$ for every $n\in\mathbb{N}$ and every $\alpha\in A$.
+    The second $\ell^{2}$-Betti number is $b^{(2)}_{2}(G_{x})=x$. Further, $b^{(2)}_{n}(G_{x})<\infty$ for every $n\in\mathbb{N}$ and every $x\in\mathbb{R}_{>0}$.
   4. 4.
-    The $\ell^{2}$-Euler characteristic is $\chi^{(2)}(G_{\alpha})=b_{2}^{(2)}(G_{\alpha})+\chi$.
+    The $\ell^{2}$-Euler characteristic is $\chi^{(2)}(G_{x})=b^{(2)}_{2}(G_{x})+\chi$.
 By renouncing simplicity, we can also construct groups where a single prime obstructs the finiteness of the cohomological dimension: see Theorem 6.1.
@@
 Lacunary hyperbolicity distinguishes them from the counterexamples to the general Atiyah Conjecture, since these all contain lamplighters [Gra16, Problem 3], while lacunary hyperbolic groups cannot (Proposition 4.15).
+Focusing on rational numbers, a simpler version of the construction leading to A allows to build hyperbolic groups.
+Theorem B.
+There exists an integer $\chi\geqslant 1$ such that the following holds. There exists a family of groups $(G_{x})_{x\in\mathbb{Q}_{>0}}$ with the following properties.
+  1. 1.
+    $G_{x}$ has property $(T)$ and is hyperbolic.
+  2. 2.
+    $G_{x}$ has cohomological dimension $16$ modulo $\pi$, where $\pi$ is the set of prime factors of the denominator of $x\in\mathbb{Q}_{>0}$.
+  3. 3.
+    The second $\ell^{2}$-Betti number is $b^{(2)}_{2}(G_{x})=x$.
+  4. 4.
+    The $\ell^{2}$-Euler characteristic is $\chi^{(2)}(G_{x})=b^{(2)}_{2}(G_{x})+\chi$.
+In particular, when $x\in\mathbb{N}$, the group $G_{x}$ constructed above has finite cohomological dimension over $\mathbb{Z}$, hence is torsion-free. This construction has the following consequence.
+Corollary C.
+Let $(m_{i})_{i\in\mathbb{N}}$ be a sequence of natural numbers.
+There exist property $(T)$ torsion-free hyperbolic groups $G,G_{i}$ of cohomological dimension $16$, with quotients $G\to G_{i}$ such that $G_{i}\xrightarrow{i\to\infty}G$ in the space of marked groups, and
+$b^{(2)}_{2}(G_{i})=m_{i}\text{ for all }i\in\mathbb{N}\text{ but }b^{(2)}_{2}(G)=0.$
+Therefore the second $\ell^{2}$-Betti number is far from being semi-continuous in the space of marked groups, even with the strong finiteness properties that are enjoyed by torsion-free hyperbolic groups.
+This is in contrast with the semi-continuity of the first $\ell^{2}$-Betti number [Pic06].
 As noticed by Ioana–Tucker-Drob, it follows easily from Gaboriau’s construction [Gab00, Proposition VI.16] that there exist uncountably many pairwise non-measure equivalent groups [ITD25, Proposition 5.1].
@@
 In particular, these groups have infinite rational cohomological dimension.
-Corollary B.
+Corollary D.
 The groups from A are pairwise non-measure equivalent.
 Proof.
-Let $G_{\alpha}$ and $G_{\beta}$ be groups from the family.
-If they are measure equivalent with index $c>0$, then $b^{(2)}_{n}(G_{\alpha})=cb^{(2)}_{n}(G_{\beta})$ for all $n\in\mathbb{N}$ by Gaboriau’s theorem [Gab02], and thus $\chi^{(2)}(G_{\alpha})=c\chi^{(2)}(G_{\beta})$.
+Let $G_{x}$ and $G_{y}$ be groups from the family.
+If they are measure equivalent with index $c>0$, then $b^{(2)}_{n}(G_{x})=cb^{(2)}_{n}(G_{y})$ for all $n\in\mathbb{N}$ by Gaboriau’s theorem [Gab02], and thus $\chi^{(2)}(G_{x})=c\chi^{(2)}(G_{y})$.
 Hence
-$cb^{(2)}_{2}(G_{\beta})+\chi=b^{(2)}_{2}(G_{a})+\chi=\chi^{(2)}(G_{\alpha})=c\chi^{(2)}(G_{\beta})=cb^{(2)}_{2}(G_{\beta})+c\chi.$
-Since $\chi$ is a positive integer, this shows that $c=1$, and thus $b^{(2)}_{2}(G_{\alpha})=b^{(2)}_{2}(G_{\beta})$, so $\alpha=\beta$. ∎
-A version of the construction leading to A produces torsion-free hyperbolic groups with prescribed integer second $\ell^{2}$-Betti number (Theorem 5.1). This has the following consequence.
-Theorem C.
-Let $(n_{i})_{i\in\mathbb{N}}$ be a sequence of natural numbers.
-There exist property $(T)$ torsion-free hyperbolic groups $G,G_{i}$ of cohomological dimension $16$, with quotient maps $G\to G_{i}$ such that $G_{i}\xrightarrow{i\to\infty}G$ in the space of marked groups, and
-$b^{(2)}_{2}(G_{i})=n_{i}\text{ for all }i\in\mathbb{N}\text{ but }b^{(2)}_{2}(G)=0.$
-Therefore the second $\ell^{2}$-Betti number is far from being semi-continuous in the space of marked groups, even with the strong finiteness properties that are enjoyed by torsion-free hyperbolic groups.
-This is in contrast with the semi-continuity of the first $\ell^{2}$-Betti number [Pic06].
+$cb^{(2)}_{2}(G_{y})+\chi=b^{(2)}_{2}(G_{x})+\chi=\chi^{(2)}(G_{x})=c\chi^{(2)}(G_{y})=cb^{(2)}_{2}(G_{y})+c\chi.$
+Since $\chi$ is a positive integer, this shows that $c=1$, and thus $b^{(2)}_{2}(G_{x})=b^{(2)}_{2}(G_{y})$, so $x=y$. ∎
+The same argument gives, to the best of our knowledge, the first infinite family of measurably diverse hyperbolic groups with property $(T)$.
+Corollary E.
+The groups from B are pairwise non-measure equivalent. ∎
 Our computations of $\ell^{2}$-Betti numbers exploit the work of Petrosyan–Sun on group-theoretic Dehn filling [PS24a].
@@
 Under this condition, building on an observation from [FF25], we will see that the excision theorem from [PS24a] is sufficient to give exact computations of the second $\ell^{2}$-Betti number of iterated group-theoretic Dehn fillings.
+Let us mention that this construction could also be done in the setting of relatively hyperbolic groups, starting from non-cocompact lattices in the octonionic hyperbolic plane, thanks to the recent work [LNP25]; this would produce groups of cohomological dimension $15$ instead.
 All of the constructions mentioned above, both in this paper and in others, are non-elementary to various extents.
 Because of this, we feel it is worth presenting possibly the most elementary construction of finitely generated groups achieving uncountably many values of $\ell^{2}$-Betti numbers.
-Theorem D.
+Theorem F.
 Let $p$ be a prime.
@@
 FFF is supported by the Herchel Smith Postdoctoral Fellowship Fund.
-The authors would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Operators, Graphs, Groups, where work on this paper was undertaken.
+The authors thank Antonio López Neumann and Bin Sun for comments on a previous version of this paper.
+They thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Operators, Graphs, Groups, where work on this paper was undertaken.
 This work was supported by EPSRC grant EP/Z000580/1.
@@
 Proposition 2.2 ([Bie81, Proposition 4.8 on p. 61]).
-Let $G_{1}\to G_{2}\to\cdots$ be a directed sequence of groups with colimit $G_{\infty}$, and let $M$ be an $R[G_{\infty}]$-module. Then there is an isomorphism
+Let $G_{0}\to G_{1}\to\cdots$ be a directed sequence of groups with colimit $G_{\infty}$, and let $M$ be an $R[G_{\infty}]$-module. Then there is an isomorphism
 $H_{n}(G_{\infty};M)\cong\operatorname{colim}H_{n}(G_{i};M).$
@@
 This is of course not true in cohomology, but it does work under some additional assumptions.
-We point out that the following is only used to compute cohomological dimensions of the groups from A and D.
+We point out that the following is only used to compute cohomological dimensions of the groups from Theorems A and F.
 If the reader is happy with the standard estimate of cohomological dimension as one more than the homological dimension [Bie81, Theorem 4.6 on p.
@@
 Proposition 2.3.
-Let $G_{1}\to G_{2}\to\cdots$ be a directed sequence of groups with colimit $G_{\infty}$, and let $M$ be an $R[G_{\infty}]$-module.
+Let $G_{0}\to G_{1}\to\cdots$ be a directed sequence of groups with colimit $G_{\infty}$, and let $M$ be an $R[G_{\infty}]$-module.
 Suppose that the pullback $H^{n}(G_{i+1};M)\to H^{n}(G_{i};M)$ is surjective for all $i\in\mathbb{N}$.
@@
 We will replace this with the stronger and easier to check condition that each map $D_{i+1}\to D_{i}$ is surjective.
-Let $X_{1}$ be a model of the classifying space $BG_{1}$111also denoted by $K(G_{1},1)$..
+Let $X_{0}$ be a model of the classifying space $BG_{0}$111also denoted by $K(G_{0},1)$..
 Using mapping cones, we may inductively define $X_{i+1}$ as a model of $BG_{i+1}$ such that the map $X_{i}\to X_{i+1}$ is an inclusion of CW-complexes.
@@
 Corollary 2.4.
-Let $G_{1}\to G_{2}\to\cdots$ be a directed sequence of groups with colimit $G_{\infty}$. Suppose that the following hold for all $i\in\mathbb{N}$.
+Let $G_{0}\to G_{1}\to\cdots$ be a directed sequence of groups with colimit $G_{\infty}$. Suppose that the following hold for all $i\in\mathbb{N}$.
   • For every $R[G_{\infty}]$-module $M$ the pullback $H^{n}(G_{i+1};M)\to H^{n}(G_{i};M)$ is surjective.
   • $cd_{R}(G_{i})=n$;
-  • $H^{n}(G_{1};R)\neq 0$ and the pullback $H^{n-1}(G_{i+1};R)\to H^{n-1}(G_{i};R)$ is surjective.
+  • $H^{n}(G_{0};R)\neq 0$ and the pullback $H^{n-1}(G_{i+1};R)\to H^{n-1}(G_{i};R)$ is surjective.
 Then $cd_{R}(G_{\infty})=n$.
@@
 The embedding $\mathcal{N}(\pi)\hookrightarrow V_{\pi}$ of from Eq. 1 is rank dense:
-One calls a $\mathcal{N}(\pi)$-linear homomorphism $f\colon M\to N$ of $\mathcal{N}(\pi)$-modules rank dense if for every $x\in N$ there is an increasing sequence of projections $p_{i}\in\mathcal{N}(\pi),i\in\mathbb{N}$, converging to $\mathrm{id}$ such that $p_{i}x\in N$.
+One calls a $\mathcal{N}(\pi)$-linear homomorphism $f\colon M\to N$ of $\mathcal{N}(\pi)$-modules rank dense if for every $x\in N$ there is an increasing sequence of projections $p_{i}\in\mathcal{N}(\pi),i\in\mathbb{N}$, converging to $\mathrm{id}$ such that $p_{i}x\in f(M)$.
 By [KPV15, Lemma A.5] the embedding $\mathcal{N}(\pi)\hookrightarrow V_{\pi}$ is rank dense because it is dense.
@@
 Let $G$ be a cocompact lattice in the isometry group of the octonionic hyperbolic plane. Then $G$ has properties $(T_{3})$ and $T_{2}]$.
-In the proofs of Theorems A and C, we will actually only need that $G$ has property $(T_{2})$: see Remark 5.3.
+Cocompact lattices in the isometry group of the octonionic hyperbolic plane exist [BHC62], and they can be chosen torsion-free up to passing to a finite index subgroup [Sel60].
+In the proofs of our main results, we will actually only need that $G$ has property $(T_{2})$: see Remark 5.3.
 Proof.
@@
 Theorem 4.2 ([PS24a, Theorem A (ii)], see also [PS24b, Corollary 4.5]).
-Let $(G,P,N)$ be a Cohen–Lyndon triple. Then for all $\mathbb{Z}[\bar{G}]$-modules $M$ and all $n\in\mathbb{N}$, the quotient maps $G\to\bar{G}$ and $P\to\bar{P}$ induce isomorphisms
+Let $(G,P,N)$ be a Cohen–Lyndon triple. Then for all $\mathbb{Z}[\bar{G}]$-modules $M$ and all $n\in\mathbb{N}$, the quotients $G\to\bar{G}$ and $P\to\bar{P}$ induce isomorphisms
 Hn(G,P;M)Hn(G¯,P¯;M) and $H_{n}(G,P;M)\cong H_{n}(\bar{G},\bar{P};M)\quad\text${ and $\quad$ $\cong H^{n}(\bar{G},\bar{P};M)$.
-We will only be concerned with an especially well-behaved case.
+We will mostly be concerned with an especially well-behaved case.
 Corollary 4.3.
@@
 Let $R$ be a ring where $[P:N]$ is invertible.
-Then for all $R[\bar{G}]$-modules $M$ and all $n\geqslant 3$, the quotient map $G\to\bar{G}$ induces isomorphisms
+Then for all $R[\bar{G}]$-modules $M$ and all $n\geqslant 3$, the quotient $G\to\bar{G}$ induces isomorphisms
 Hn(G;M)Hn(G¯;M) and $H_{n}(G;M)\cong H_{n}(\bar{G};M)\quad\text${ and $\quad$ $\cong H^{n}(\bar{G};M)$.
@@
 ∎
+A useful feature of the Cohen–Lyndon property is that it is transitive. The following is a generalisation of [CIOS26, Lemma 4.22], which deals with the case in which $L=N$.
+Lemma 4.4.
+Let $N<L<P<G$ be groups such that $(G,P,\left<\!\left<N\right>\!\right>_{P})$ and $(P,L,N)$ are Cohen–Lyndon triples (in particular $N$ is normal in $L$). Then $(G,L,N)$ is a Cohen–Lyndon triple.
+Proof.
+Let $T$ be a set of left transversals of $G/P\left<\!\left<N\right>\!\right>_{G}$ such that $\left<\!\left<N\right>\!\right>_{G}=*_{t\in T}t\left<\!\left<N\right>\!\right>_{P}t^{-1}$, and let $S$ be a set of left transversals of $P/L\left<\!\left<N\right>\!\right>_{P}$ such that $\left<\!\left<N\right>\!\right>_{P}=*_{s\in S}sNs^{-1}$.
+Hence
+$\left<\!\left<N\right>\!\right>_{G}=*_{t\in T}t\left<\!\left<N\right>\!\right>_{P}t^{-1}=*_{t\in T,s\in S}tsN(ts)^{-1},$
+so it remains to show that $TS$ is a set of left transversals for $G/L\left<\!\left<N\right>\!\right>_{G}$.
+For $g\in G$, there exist $t\in T$ and $p\in P$ such that $g\in tp\left<\!\left<N\right>\!\right>_{G}$. Further, there exists $s\in S$ such that $p\in sL\left<\!\left<N\right>\!\right>_{P}$. Hence
+$g\in tp\left<\!\left<N\right>\!\right>_{G}\subset t(sL\left<\!\left<N\right>\!\right>_{P})\left<\!\left<N\right>\!\right>_{G}=tsL\left<\!\left<N\right>\!\right>_{G}.$
+Suppose that there exist $t_{i}\in T,s_{i}\in S$ such that $t_{1}s_{1}\in t_{2}s_{2}L\left<\!\left<N\right>\!\right>_{G}$. Then
+$t_{2}^{-1}t_{1}\in s_{2}L\left<\!\left<N\right>\!\right>_{G}s_{1}^{-1}=s_{2}Ls_{1}^{-1}\left<\!\left<N\right>\!\right>_{G}\subset P\left<\!\left<N\right>\!\right>_{G}.$
+Because $T$ is a set of left transversals of $G/P\left<\!\left<N\right>\!\right>_{G}$, we have $t_{1}=t_{2}$, so $s_{2}^{-1}s_{1}\in L\left<\!\left<N\right>\!\right>_{G}$.
+Let $l\in L$ be such that $l^{-1}s_{2}^{-1}s_{1}\in\left<\!\left<N\right>\!\right>_{G}$.
+Now $l^{-1}s_{2}^{-1}s_{1}\in P$, and $\left<\!\left<N\right>\!\right>_{G}\cap P=\left<\!\left<N\right>\!\right>_{P}$ because $(G,P,\left<\!\left<N\right>\!\right>_{P})$ is a Cohen–Lyndon triple [Sun20, Proposition 6.1(a)], so $l^{-1}s_{2}^{-1}s_{1}\in\left<\!\left<N\right>\!\right>_{P}$, that is $s_{1}\in s_{2}L\left<\!\left<N\right>\!\right>_{P}$.
+Because $S$ is a set of left transversals for $P/L\left<\!\left<N\right>\!\right>_{P}$, we have $s_{1}=s_{2}$ and conclude.
+∎
 4.2 (Relatively) hyperbolic groups
@@
 Theorem 4.6 ([Osi07]).
-Let $G$ be hyperbolic relative to $\{P_{\lambda}\}_{\lambda\in\Lambda}$. Let $B\subset G$ be a finite subset. Then for every sufficiently deep normal subgroups $N_{\lambda}<P_{\lambda}$, it holds
+[…]
   1. 1.
-[…]
+    The group $\bar{G}$ is hyperbolic relative to $\bar{P}$;
   2. 2.
-[…]
-  3. 3.
-[…]
-We will mostly be concerned with a very particular example of a relatively hyperbolic pair.
-Let $G$ be a hyperbolic group, and let $g\in G$ be an element of infinite order.
-Then $g$ is contained in a unique maximal virtually cyclic subgroup denoted $E_{G}(g)$.
+    The quotient $G\to\bar{G}$ is injective on $B$ and induces embedding $\bar{P}\to\bar{G}$.
+We will mostly be concerned with a particular example of a relatively hyperbolic pair. Recall that a subgroup $P<G$ is almost malnormal if $gPg^{-1}\cap P$ is finite for all $g\in G\setminus P$.
 Theorem 4.7 ([Bow12, Theorem 7.11]).
-Let $G$ be a hyperbolic group, and let $g\in G$ be an element of infinite order. Then $G$ is hyperbolic relative to $E_{G}(g)$.
-Putting these results together we obtain the following, which will serve as the inductive step in our constructions.
+Let $G$ be a non-elementary hyperbolic group, and let $P<G$ be quasiconvex and almost malnormal. Then $G$ is hyperbolic relative to $P$.
+The main example is the following.
+Definition 4.8.
+Let $G$ be a non-elementary hyperbolic group, and $g\in G$ an element of infinite order.
+The elementary closure of $g$ is the unique maximal virtually cyclic subgroup $E_{G}(g)<G$ containing $g$.
+We say that $g\in G$ is special if it has infinite order and $E_{G}(g)=\langle g\rangle$.
+Starting from a quasiconvex almost malnormal subgroup, we can build a new one.
+The following notion is central in much of the literature on small cancellation theory on groups with hyperbolic features [Ols93, Osi10, Hul16].
+Definition 4.9.
+Let $G$ be a non-elementary hyperbolic group. A subgroup $H<G$ is suitable if it is non-elementary and does not normalise any non-trivial finite subgroup of $G$.
+Once again there is a main example we will care about.
+Lemma 4.10 ([Osi25, Lemma 2.4]).
+Let $G$ be a non-elementary hyperbolic group with no non-trivial finite normal subgroups. Then every non-trivial normal subgroup of $G$ is suitable.
+The following, which is the consequence of the more general [AH21, Theorem 6.2], says that we can extend quasiconvex almost malnormal subgroups via elements of suitable subgroups.
+Theorem 4.11 ([AH21]).
+Let $G$ be a non-elementary hyperbolic group with no non-trivial finite normal subgroups, and let $P<G$ be an infinite-index quasiconvex almost malnormal subgroup.
+Let $H<G$ be a suitable subgroup.
+Then there exists an element $h\in H$ such that $\langle P,h\rangle\cong P*\langle h\rangle$ is an infinite-index quasiconvex almost malnormal subgroup.
+We can now use these results to prove the two main technical propositions of this section, which will be used as the induction step in our constructions.
+The first is about Dehn filling with respect to a large power of an arbitrary element.
 Proposition 4.12.
@@
   3. 3.
-    The quotient map $G\to\bar{G}$ is injective when restricted to $B$, and induces an embedding $\bar{P}\hookrightarrow\bar{G}$.
+    The quotient $G\to\bar{G}$ is injective on $B$, and induces an embedding $\bar{P}\hookrightarrow\bar{G}$.
 Proof.
-By Theorem 4.6, $G$ is hyperbolic relative to $E_{G}(g)$.
+By Theorem 4.7, $G$ is hyperbolic relative to $P$.
 The Cohen–Lyndon property follows from Theorem 4.5; the hyperbolicity follows from Theorem 4.6, since a group that is hyperbolic relative to a finite subgroup is hyperbolic; the injectivity follows again from Theorem 4.6.
 It remains to prove that $\bar{G}$ is non-elementary and without finite normal subgroups.
-To this end, we find an additional element of infinite order $h\in G$ such that $G$ is hyperbolic relative to $\{E_{G}(g),E_{G}(h)\}$ [Osi07, Lemmas 5.5 and 5.6], see also [Osi06a, Corollaries 1.7 and 4.5].
-Consider the normal subgroups $N<E_{G}(g)$ and $\{1\}<E_{G}(h)$, so that the quotient $G\to\bar{G}$ fits into the framework of Theorem 4.5.
-In particular, $\bar{G}$ is hyperbolic relative to the image of $E_{G}(h)$, which maps injectively and is almost malnormal in $\bar{G}$ by [Osi06b, Theorem 1.4].
-If $\bar{G}$ had a non-trivial finite normal subgroup, then a sufficiently high power of $h$ would centralise it, contradicting malnormality.
-Similarly, if $\bar{G}$ were virtually cyclic, then a sufficiently high power of $h$ would generate a normal subgroup of $\bar{G}$, contradicting malnormality again.
+To this end, we use Theorem 4.11 to find an additional element $h\in G$ such that $G$ is hyperbolic relative to $P\coloneqq E_{G}(g)*\langle h\rangle$.
+We can now see $G/\left<\!\left<N\right>\!\right>_{G}$ as a Dehn filling of $(G,P)$ with respect to the sufficiently deep normal subgroup $\left<\!\left<N\right>\!\right>_{P}<P$.
+In particular, $\bar{G}$ is hyperbolic relative to $P/\left<\!\left<N\right>\!\right>_{P}\cong\bar{P}*\langle h\rangle$.
+Since this is not virtually cyclic and has no non-trivial finite normal subgroups, the same is true for $G$, by [DGO17, Theorem 6.14(b)].
 ∎
-The situation when $E_{G}(g)=\langle g\rangle$ is especially convenient. Following [OT13], we call such elements special.
-Proposition 4.8 ([OT13, Proposition 3.4]).
-[…]
-In the cases above, to obtain desireable properties we needed to pass to a sufficiently deep normal subgroup of the elementary closure of an element.
-Let us highlight one situation in which this is not necessary.
-Lemma 4.9.
+The second result is about Dehn filling with respect to a prescribed power of a well-chosen element.
+Proposition 4.13.
 […]
   1. 1.
-    The triple $(G,\langle g\rangle,\langle g\rangle)$ is Cohen–Lyndon;
+    The triple $(G,\langle g\rangle,\langle g^{k}\rangle)$ is Cohen–Lyndon;
   2. 2.
-    The group $\bar{G}$ is torsion-free non-elementary hyperbolic;
+    The group $\bar{G}$ is non-elementary hyperbolic with no non-trivial finite normal subgroups;
   3. 3.
-    The quotient map $G\to\bar{G}$ is injective when restricted to $B$.
+    The quotient $G\to\bar{G}$ is injective on $B$, and the image of $g$ has order $k$.
 Proof.
-This obtained in the course of the proof of [FFS25, Theorem 3.1], the second and third statement are consequences of the theorem, the first one is [FFS25, Lemma 3.5]. ∎
+We first find a special element $y\in xH$.
+By [AMO07, Lemma 3.8], the suitable subgroup $H$ contains a special element $h$.
+If $x\in H$, then we can set $y=h\in xH$; otherwise [OT13, Lemma 3.2] ensures that $x\langle h\rangle$ contains a special element $y$.
+So $E_{G}(y)=\langle y\rangle$ is infinite-index, quasiconvex and almost malnormal in $G$.
+Applying Theorem 4.11 twice, we find elements $z,w\in H$ such that $P\coloneqq\langle y,z,w\rangle$ is a free group of rank $3$ that is infinite-index, quasiconvex and almost malnormal.
+By Theorem 4.7, $G$ is hyperbolic relative to $P$.
+We choose
+$g\coloneqq yzwz^{2}w^{2}\cdots z^{n}w^{n},$
+where $n$ is large enough such that $g$ is a $C^{\prime}(\lambda)$-small cancellation word in $P$, for some small parameter $\lambda>0$.
+By classical small cancellation theory (see e.g.
+[LS01, Chapter V]), the normal subgroup $\left<\!\left<g\right>\!\right>_{P}<P$ is sufficiently deep, hence so is $\left<\!\left<g^{k}\right>\!\right>_{P}$.
+We can thus apply Theorem 4.6 and obtain that $G/\left<\!\left<g^{k}\right>\!\right>_{G}$ is hyperbolic relative to $P/\left<\!\left<g^{k}\right>\!\right>_{P}$, and that both $B$ and $P/\left<\!\left<g^{k}\right>\!\right>_{P}$ embed into $\bar{G}$.
+If $k=1$, then $P/\left<\!\left<g^{k}\right>\!\right>_{P}=P/\left<\!\left<g\right>\!\right>_{P}$ is a classical small cancellation group, hence it is hyperbolic [Gro87].
+If $k>1$, then $P/\left<\!\left<g^{k}\right>\!\right>_{P}$ is a one-relator group with torsion, hence Newman’s spelling theorem [New68] implies that it is hyperbolic, and that the image of $g$ has order $k$.
+In both cases, $\bar{G}$ is hyperbolic relative to a non-elementary hyperbolic group, hence it is itself non-elementary hyperbolic.
+Since $P/\left<\!\left<g^{k}\right>\!\right>_{P}$ also has no non-trivial finite normal subgroups in both cases, the same is true of $\bar{G}$ by [DGO17, Theorem 6.14(b)].
+Finally, in both cases $g$ has order $k$ in $P/\left<\!\left<g^{k}\right>\!\right>_{P}$, which embeds in $\bar{G}$, so $g$ has order $k$ in $\bar{G}$ as well.
+It remains to show the first item.
+Since $\left<\!\left<g^{k}\right>\!\right>_{P}<P$ is sufficiently deep, by Theorem 4.5 the triple $(G,P,\left<\!\left<g^{k}\right>\!\right>_{P})$ is Cohen–Lyndon.
+Moreover by the Cohen–Lyndon theorem on one-relator groups [CL63], the triple $(P,\langle g\rangle,\langle g^{k}\rangle)$ is Cohen–Lyndon.
+So Lemma 4.4 implies that $(G,\langle g\rangle,\langle g^{k}\rangle)$ is Cohen–Lyndon.
+∎
 4.3 Lacunary hyperbolic groups
@@
 Theorem 4.14 ([OOS09, Theorem 1.1]).
-Let $G_{1}\to G_{2}\to\cdots$ be a directed sequence of epimorphisms of hyperbolic groups, with colimit $G_{\infty}$.
-Let $S_{1}$ be a finite generating set for $G_{1}$, and let $S_{i}$ be its pushforward to $G_{i}$; let $\delta_{i}$ be the hyperbolicity constant of $\mathrm{Cay}(G_{i},S_{i})$.
+Let $G_{0}\to G_{1}\to\cdots$ be a directed sequence of epimorphisms of hyperbolic groups, with colimit $G_{\infty}$.
+Let $S_{0}$ be a finite generating set for $G_{0}$, and let $S_{i}$ be its pushforward to $G_{i}$; let $\delta_{i}$ be the hyperbolicity constant of $\mathrm{Cay}(G_{i},S_{i})$.
 For each $i\in\mathbb{N}$, let $\rho_{i}$ be the radius of the largest ball centred at the identity on which the quotient $G_{i}\to G_{i+1}$ is injective.
@@
 ∎
-5 Failure of semi-continuity
-In this section we prove C. We start with a more precise statement.
+5 Rationals and hyperbolic groups
+In this section we prove B and C.
+5.1 Realising natural numbers
+We start with the torsion-free case, which is simpler, and sufficient for C.
 Theorem 5.1.
 Let $G$ be a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane, and let $B\subset G$ be a finite subset.
-Then for every $n\in\mathbb{N}$, there exists a torsion-free hyperbolic quotient $G\to G_{n}$, injective on $B$, of cohomological dimension $16$, such that for every unitary $G_{n}$-representation $V$, the second homology is $H_{2}(G_{n};V)\cong V^{n}$.
-Cocompact lattices in the isometry group of the octonionic hyperbolic plane exist [BHC62], and they can be chosen torsion-free up to passing to a finite index subgroup [Sel60].
+Then for every $m\in\mathbb{N}$, there exists a torsion-free hyperbolic quotient $G\to G_{m}$, injective on $B$, of cohomological dimension $16$, such that for every unitary $G_{m}$-representation $V$, the second homology is $H_{2}(G_{m};V)\cong V^{m}$.
 Proof.
 The group $G$ is torsion-free hyperbolic of cohomological dimension $16$, and by Theorem 3.10 it has property $T_{2}]$.
-We construct $G_{n}$ by induction on $n$, starting with $G_{0}=G$, which works by Proposition 3.8.
-Suppose by induction that $G_{n-1}$ has been constructed.
-Lemma 4.9 gives an element $g\in G_{n-1}$ and we define $G_{n}$ to be the corresponding group-theoretic Dehn filling.
-Lemma 4.9 gives injectivity on $B$ and the fact that $G_{n}$ is torsion-free hyperbolic.
-The first part of Corollary 4.3 implies that the cohomological dimension of $G_{n}$ is $16$.
-Moreover, since $G_{n-1}$ has property $(T)$, applying Proposition 3.8 and the second part of Corollary 4.3, we obtain, for every unitary $G_{n}$-representation $V$, a short exact sequence
-$0\to H_{2}(G_{n-1};V)\to H_{2}(G_{n};V)\to V\to 0.$
-By the induction hypothesis the first term is $V^{n-1}$. Therefore $H_{2}(G_{n};V)\cong V^{n}$, which proves the claim and hence the theorem. ∎
+We construct $G_{m}$ by induction on $n$, starting with $G_{0}=G$, which works by Proposition 3.8.
+Suppose by induction that $G_{m}$ has been constructed.
+Apply Proposition 4.13 with the finite set $B$ (and any $x,H$), let $g$ be the resulting element and $G_{m+1}$ the Dehn filling quotient.
+Hence $G_{m+1}$ is non-elementary hyperbolic, the quotient $G\to G_{m+1}$ is injective on $B$, and $(G_{m},\langle g\rangle,\langle g\rangle)$ is a Cohen–Lyndon triple.
+Now $G$ is the fundamental group of a closed $16$-dimensional aspherical manifold, so it has cohomological dimension $16$, and $H^{16}(G;\mathbb{Z})\cong\mathbb{Z}$.
+The first part of Corollary 4.3 therefore implies that the cohomological dimension of $G_{m+1}$ is $16$, hence in particular that $G_{m+1}$ is torsion-free.
+Moreover, since $G_{m}$ has property $(T)$, applying Proposition 3.8 and the second part of Corollary 4.3, we obtain, for every unitary $G_{m+1}$-representation $V$, a short exact sequence
+$0\to H_{2}(G_{m};V)\to H_{2}(G_{m+1};V)\to V\to 0.$
+By the induction hypothesis the first term is $V^{m}$. Therefore $H_{2}(G_{m+1};V)\cong V^{m+1}$, which proves the claim and hence the theorem. ∎
 Remark 5.2.
-The groups $G_{n}$ of course fail to have property $(T_{2})$, despite being quotients of the $(T_{2})$ group $G$, as remarked in [FF25].
-Still, the second homology (and similarly the second cohomology) functor enjoys some strong form of rigidity: in a way, its output can be read directly from its input.
+The groups $G_{m}$ of course fail to have property $(T_{2})$, despite being quotients of the $(T_{2})$ group $G$, as remarked in [FF25].
+Still, the second homology (and similarly the second cohomology) functor enjoys some strong form of rigidity: in a way, its output can be read directly from its input, and the same will be true in the next constructions.
 It would be interesting to explore whether a (co)homological rigidity of this form characterises quotients of $(T_{2})$ groups.
@@
 We have $b^{(2)}_{2}(G)=0$ by Corollary 3.9.
-For each $i\in\mathbb{N}$, we let $B_{i}$ be the ball of radius $i$ in $G$, for a fixed choice of finite generating set, and $n_{i}\in\mathbb{N}$ as in the statement of C.
+For each $i\in\mathbb{N}$, we let $B_{i}$ be the ball of radius $i$ in $G$, for a fixed choice of finite generating set, and $m_{i}\in\mathbb{N}$ as in the statement of C.
 Then Theorem 5.1 builds a group $G_{i}$, which by Proposition 3.6 has
-$b^{(2)}_{2}(G_{i})=\dim_{\mathcal{N}(G_{i})}H_{2}(G_{i};\ell^{2}(G_{i}))=\dim_{\mathcal{N}(G_{i})}\ell^{2}(G_{i})^{n_{i}}=n_{i}.$
+$b^{(2)}_{2}(G_{i})=\dim_{\mathcal{N}(G_{i})}H_{2}(G_{i};\ell^{2}(G_{i}))=\dim_{\mathcal{N}(G_{i})}\ell^{2}(G_{i})^{m_{i}}=m_{i}.$
 Since $G\to G_{i}$ is injective on larger and larger balls, $G_{i}\xrightarrow{i\to\infty}G$ in the space of marked groups.
@@
 In the proof of C, we applied Theorem 5.1 only with $V=\ell^{2}(G_{i})$, which has no non-zero invariant vectors, so property $(T_{2})$ would have been sufficient.
-Similarly, the computation of second $\ell^{2}$-Betti numbers in the next section could be obtained only using property $(T_{2})$.
-6 Proof of the main theorem
+Similarly, the computation of second $\ell^{2}$-Betti numbers in the rest of this section and in the next could be obtained only using property $(T_{2})$.
+5.2 Realising rational numbers
+Next we consider arbitrary rational numbers, completing the proof of B.
+Theorem 5.4 (corresponding to B).
+Let $x\in\mathbb{Q}_{\geqslant 0}$ and let $\pi$ be the set of prime factors of the denominator of $x$.
+Then there exists a hyperbolic group $G$ with property $(T)$ and cohomological dimension $16$ modulo $\pi$, such that $b^{(2)}_{2}(G)=x$.
+Proof.
+Let $G_{0}$ be a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane, which has Euler characteristic $\chi\geqslant 1$ [Lüc02, Corollary 5.16(1) on p.
+231].
+If $x=0$, then we are done.
+Otherwise let $x=\frac{m}{k}$ be a reduced fraction, where $m,k\in\mathbb{N}$.
+We will build by induction a sequence of epimorphisms of hyperbolic groups $G_{0}\to G_{1}\to\cdots\to G_{m}$ and elements $g_{i}\in G_{i}$, such that:
+  1. 1.
+    $G_{i}$ is non-elementary hyperbolic with no non-trivial finite normal subgroups, and cohomological dimension $16$ modulo $\pi$.
+  2. 2.
+    The element $g_{i}\in G_{i}$ has infinite order, and for all $j>i$ its image in $G_{j}$ has order $k$.
+  3. 3.
+    Each quotient $G_{i}\to G_{i+1}$ induces isomorphisms in homology and cohomology with coefficients in $R[G_{i+1}]$-modules and degree at least $3$, where $R$ is any ring where every prime from $\pi$ is invertible.
+  4. 4.
+    For every unitary $G_{i}$-representation $V$, the second homology is
+    $H_{2}(G_{i};V)=\bigoplus_{j<i}V_{g_{j}}.$
+The group $G=G_{m}$ is then non-elementary hyperbolic, with no non-trivial finite normal subgroups and cohomological dimension $16$ modulo $\pi$.
+It is a quotient of $G_{0}$ hence it has property $(T)$.
+Each $g_{i}:i<m$ has order $k$ in $G$, so by Propositions 3.6 and 3.11:
+$\dim_{\mathcal{N}(G)}\ell^{2}(G)_{g_{i}}$ $=\dim_{\mathcal{N}(G)}H_{0}(\langle g_{i}\rangle,\ell^{2}(G))=\dim_{\mathcal{N}(G)}H_{0}(\langle g_{i}\rangle,\mathcal{N}(G))$
+$=\dim_{\mathcal{N}(\langle g_{i}\rangle)}H_{0}(\langle g_{i}\rangle,\mathcal{N}(\langle g_{i}\rangle))=b^{(2)}_{0}(\langle g_{i}\rangle)=\frac{1}{k}.$
+By Proposition 3.6 and additivity we therefore have
+$b^{(2)}_{2}(G)=\dim_{\mathcal{N}(G)}H_{2}(G;\ell^{2}(G))=\sum\limits_{0\leqslant i<m}\frac{1}{k}=x.$
+By Item 3, for all $n\geqslant 3$ we have
+$b^{(2)}_{n}(G_{m})=\dim_{\mathcal{N}(G_{x})}H_{n}(G_{m};\mathcal{N}(G_{m}))=\dim_{\mathcal{N}(G_{m})}H_{n}(G_{0};\mathcal{N}(G_{m}))=b^{(2)}_{n}(G_{0}\to G_{x}).$
+Since $G_{m}$ has property $(T)$, $b^{(2)}_{1}(G_{m})=0$ by [BV97, Corollary 6]. Hence
+$\chi^{(2)}(G_{m})$ $=\sum\limits_{n\in\mathbb{N}}(-1)^{n}b^{(2)}_{n}(G_{m})=b^{(2)}_{2}(G_{m})+\sum\limits_{n\geqslant 3}(-1)^{n}b^{(2)}_{n}(G_{m})$
+$=b^{(2)}_{2}(G_{m})+\sum\limits_{n\geqslant 3}(-1)^{n}b^{(2)}_{n}(G_{0}\to G_{m})=b^{(2)}_{2}(G_{m})+\chi,$
+where the last equality follows from Corollary 3.9.
+It remains to construct the sequence.
+Suppose by induction that $G_{i}$ has been constructed, for some $0\leqslant i<n$.
+Apply Proposition 4.13 with $k$ (and any $x,H$), let $g_{i}$ be the resulting element and $G_{i+1}$ the Dehn filling quotient; we also choose $B$ to be sufficiently large such that for all $j<i$, the element $g_{j}$ still has order $k$ in $G_{i+1}$.
+Hence $G_{i+1}$ is non-elementary hyperbolic with no non-trivial finite normal subgroup, and $(G_{i},\langle g_{i}\rangle,\langle g_{i}^{k}\rangle)$ is a Cohen–Lyndon triple.
+Induction and the first part of Corollary 4.3 imply that $G_{i+1}$ has cohomological dimension $16$ modulo $\pi$.
+Moreover, the second part and property $(T)$ give, for every unitary $G_{i+1}$-representation $V$, a short exact sequence
+$0\to H_{2}(G_{i};V)\to H_{2}(G_{i+1};V)\to V_{g_{i}}\to 0.$
+This concludes the construction, hence the proof. ∎
+6 Reals and simple groups
 In this section we prove A. Because the proof has many moving parts, we start with a simpler version, that only involves one prime, and does not give simplicity.
 Theorem 6.1.
-Let $p$ be a prime. There exists an uncountable family of groups $(G_{\alpha})_{\alpha\in A}$ and an integer $\chi\geq 1$ with the following properties.
+Let $p$ be a prime. There exists a family of groups $(G_{x})_{x\in\mathbb{R}_{>0}}$ and an integer $\chi\geq 1$ with the following properties.
   1. 1.
-    $G_{\alpha}$ has property $(T)$ and is lacunary hyperbolic.
+    $G_{x}$ is simple, has property $(T)$, and is lacunary hyperbolic.
   2. 2.
-    $G_{\alpha}$ has cohomological dimension $16$ modulo $p$.
+    $G_{x}$ has cohomological dimension $16$ modulo $p$.
   3. 3.
-    The second $\ell^{2}$-Betti numbers $b^{(2)}_{2}(G_{\alpha})$ are pairwise distinct.
+    The second $\ell^{2}$-Betti number is $b^{(2)}_{2}(G_{x})=x$.
   4. 4.
-    The $\ell^{2}$-Euler characteristic is $\chi^{(2)}(G_{\alpha})=b_{2}^{(2)}(G_{\alpha})+\chi$.
+    The $\ell^{2}$-Euler characteristic is $\chi^{(2)}(G_{x})=b^{(2)}_{2}(G_{x})+\chi$.
 Proof.
-We choose as an index set $A\coloneqq\{2,3\}^{\mathbb{N}}$, and fix an element $\alpha\in A$.
-The group $G_{\alpha}$ will arise as the colimit of a direct sequence of epimorphisms of hyperbolic groups $G_{1}\to G_{2}\to\cdots$, where $G_{1}$ is a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane, which has Euler characteristic $\chi\geqslant 1$ [Lüc02, Corollary 5.16(1) on p.
+We fix a positive real $x\in\mathbb{R}_{>0}$ and choose a base-$p$ representation
+$\sum\limits_{i\geqslant 1}\frac{1}{p^{k_{i}}},$
+for some $k_{i}\in\mathbb{N}$.
+To treat all numbers in a unified way, if $x$ is $p$-adic, we still choose a periodic representation.
+The group $G_{x}$ will arise as the colimit of a direct sequence of epimorphisms of hyperbolic groups $G_{0}\to G_{1}\to\cdots$, where $G_{0}$ is a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane, which has Euler characteristic $\chi\geqslant 1$ [Lüc02, Corollary 5.16(1) on p.
 231].
-We fix a finite generating set $S_{1}$ for $G_{1}$ and let $S_{i}$ be the pushforward to $G_{i}$.
+We fix a finite generating set $S_{0}$ for $G_{0}$ and let $S_{i}$ be the pushforward to $G_{i}$.
 We also specify elements $g_{i}\in G_{i}$, and will prove the following additional properties.
@@
   4. 4.
-    The element $g_{i}\in G_{i}$ has infinite order, and for all $j>i$ its image in $G_{j}$ has order $p^{a_{i}}$, where $a_{i}\in\mathbb{N}$ is a sequence with the property that $a_{i+1}-a_{i}-1$ is a positive power of $\alpha(i)\in\{2,3\}$.
+    The element $g_{i}\in G_{i}$ has infinite order, and for all $j>i$ its image in $G_{j}$ has order $p^{k_{i}}$.
   5. 5.
@@
     $H_{2}(G_{i};V)=\bigoplus_{j<i}V_{g_{j}}.$
-Let us first prove that the colimit $G_{\alpha}$ has the desired properties.
-The group $G_{1}$ has property $(T)$ by Theorem 3.10, hence so does $G_{\alpha}$, being its quotient.
-Item 2 implies that $G_{\alpha}$ is lacunary hyperbolic, by Theorem 4.10.
-Now $G_{1}$ is the fundamental group of a closed $16$-dimensional aspherical manifold, so it has cohomological dimension $16$, and $H^{16}(G_{1};R)\cong R$ for every ring $R$.
-Therefore Item 3 and Corollary 2.4 imply that $G_{\alpha}$ has cohomological dimension $16$ modulo $p$.
-Next, we compute $b^{(2)}_{2}(G_{\alpha})$. By Proposition 2.2, we have
-$H_{2}(G_{\alpha};\ell^{2}(G_{\alpha}))\cong\bigoplus_{i\geqslant 1}\ell^{2}(G_{\alpha})_{g_{i}}.$
-Each $g_{i}$ has order $p^{a_{i}}$ in $G_{\alpha}$, hence by Proposition 3.6 and Proposition 3.11:
-$\dim_{\mathcal{N}(G_{\alpha})}\ell^{2}(G_{\alpha})_{g_{i}}$ $=\dim_{\mathcal{N}(G_{\alpha})}H_{0}(\langle g_{i}\rangle,\ell^{2}(G_{\alpha}))=\dim_{\mathcal{N}(G_{\alpha})}H_{0}(\langle g_{i}\rangle,\mathcal{N}(G_{\alpha}))$
-$=\dim_{\mathcal{N}(\langle g_{i}\rangle)}H_{0}(\langle g_{i}\rangle,\mathcal{N}(\langle g_{i}\rangle))=b^{(2)}_{0}(\langle g_{i}\rangle)=p^{-a_{i}}.$
+Let us first prove that the colimit $G_{x}$ has the desired properties.
+The group $G_{0}$ has property $(T)$ by Theorem 3.10, hence so does $G_{x}$, being its quotient.
+Item 2 implies that $G_{x}$ is lacunary hyperbolic, by Theorem 4.14.
+Item 3 and Corollary 2.4 imply that $G_{x}$ has cohomological dimension $16$ modulo $p$.
+Next, we compute $b^{(2)}_{2}(G_{x})$. By Proposition 2.2, we have
+$H_{2}(G_{x};\ell^{2}(G_{x}))\cong\bigoplus_{i\geqslant 1}\ell^{2}(G_{x})_{g_{i}}.$
+Each $g_{i}$ has order $p^{k_{i}}$ in $G_{x}$, hence by Propositions 3.6 and 3.11:
+$\dim_{\mathcal{N}(G_{x})}\ell^{2}(G_{x})_{g_{i}}$ $=\dim_{\mathcal{N}(G_{x})}H_{0}(\langle g_{i}\rangle,\ell^{2}(G_{x}))=\dim_{\mathcal{N}(G_{x})}H_{0}(\langle g_{i}\rangle,\mathcal{N}(G_{x}))$
+$=\dim_{\mathcal{N}(\langle g_{i}\rangle)}H_{0}(\langle g_{i}\rangle,\mathcal{N}(\langle g_{i}\rangle))=b^{(2)}_{0}(\langle g_{i}\rangle)=p^{-k_{i}}.$
 By Proposition 3.6 and Lemma 3.2, we therefore have
-$b^{(2)}_{2}(G_{\alpha})=\dim_{\mathcal{N}(G_{\alpha})}H_{2}(G_{\alpha};\ell^{2}(G_{\alpha}))=\sum\limits_{i\geqslant 1}\frac{1}{p^{a_{i}}}.$
-Reading this number in base $p$, we see a $1$ at position $a_{i}$, then a string of 0s of length $(a_{i+1}-a_{i}-1)$ before the next $1$ at position $a_{i+1}$.
-The length of this string of 0s is either a power of $2$ or a power of $3$ depending on whether $\alpha(i)=2$ or $3$.
-This shows that we can read $\alpha$ from $b^{(2)}_{2}(G_{\alpha})$.
-Finally, we compute $\chi^{(2)}(G_{\alpha})$. By Item 3 and Proposition 2.2, for all $n\geqslant 3$ we have
-$b^{(2)}_{n}(G_{\alpha})=\dim_{\mathcal{N}(G_{\alpha})}H_{n}(G_{\alpha};\mathcal{N}(G_{\alpha}))=\dim_{\mathcal{N}(G_{\alpha})}H_{n}(G_{1};\mathcal{N}(G_{\alpha}))=b^{(2)}_{n}(G_{1}\to G_{\alpha}).$
-Since $G_{\alpha}$ has property $(T)$, $b^{(2)}_{1}(G_{\alpha})=0$ by [BV97, Corollary 6]. Hence
-$\chi^{(2)}(G_{\alpha})$ $=\sum\limits_{n\in\mathbb{N}}(-1)^{n}b^{(2)}_{n}(G_{\alpha})=b^{(2)}_{2}(G_{\alpha})+\sum\limits_{n\geqslant 3}(-1)^{n}b^{(2)}_{n}(G_{\alpha})$
-$=b^{(2)}_{2}(G_{\alpha})+\sum\limits_{n\geqslant 3}(-1)^{n}b^{(2)}_{n}(G_{1}\to G_{\alpha})=b^{(2)}_{2}(G_{\alpha})+\chi,$
+$b^{(2)}_{2}(G_{x})=\dim_{\mathcal{N}(G_{x})}H_{2}(G_{x};\ell^{2}(G_{x}))=\sum\limits_{i\geqslant 1}\frac{1}{p^{k_{i}}}=x.$
+Finally, we compute $\chi^{(2)}(G_{x})$. By Item 3 and Proposition 2.2, for all $n\geqslant 3$ we have
+$b^{(2)}_{n}(G_{x})=\dim_{\mathcal{N}(G_{x})}H_{n}(G_{x};\mathcal{N}(G_{x}))=\dim_{\mathcal{N}(G_{x})}H_{n}(G_{0};\mathcal{N}(G_{x}))=b^{(2)}_{n}(G_{0}\to G_{x}).$
+Since $G_{x}$ has property $(T)$, $b^{(2)}_{1}(G_{x})=0$ by [BV97, Corollary 6]. Hence
+$\chi^{(2)}(G_{x})$ $=\sum\limits_{n\in\mathbb{N}}(-1)^{n}b^{(2)}_{n}(G_{x})=b^{(2)}_{2}(G_{x})+\sum\limits_{n\geqslant 3}(-1)^{n}b^{(2)}_{n}(G_{x})$
+$=b^{(2)}_{2}(G_{x})+\sum\limits_{n\geqslant 3}(-1)^{n}b^{(2)}_{n}(G_{0}\to G_{x})=b^{(2)}_{2}(G_{x})+\chi,$
 where the last equality follows from Corollary 3.9.
-[…]
-By Proposition 4.7 again, the triple $(G_{i},\langle g_{i}\rangle,\langle g_{i}^{k}\rangle)$ is Cohen–Lyndon.
-Because $k$ is a power of $p$, for every ring $R$ in which $p$ is invertible, we can apply Corollary 4.3.
-The first part immediately gives Item 3.
-Moreover, the second part, together with the fact that $G_{i}$ has property $(T)$, gives a short exact sequence
+It remains to construct the directed sequence $G_{0}\to G_{1}\to\cdots$ and the elements $g_{i}\in G_{i}$ so that the items above are satisfied.
+Starting with $G_{0}$, note that Item 5 holds by Proposition 3.8, because $G_{0}$ has property $T_{2}]$, and there is nothing to check for the other items, before $g_{0}$ has been chosen.
+So suppose by induction that $G_{0},\ldots,G_{i}$ and the elements $g_{0},\ldots,g_{i-1}$ have been constructed so that all items above are satisfied.
+Apply Proposition 4.13 with $k=k_{i}$, a finite set $B$ to be specified later (and any $x,H$).
+Let $g_{i}$ be the resulting element and $G_{i+1}$ the Dehn filling quotient.
+Proposition 4.13 directly gives Item 1.
+By choosing $B$ large enough, we can ensure that Item 2 is satisfied, and that the elements $g_{j}:j<i$ maintain the same order in $G_{i+1}$; this together with Proposition 4.13 gives Item 4.
+Because $(G_{i},\langle g_{i}\rangle,\langle g_{i}^{k}\rangle)$ is Cohen–Lyndon, again by Proposition 4.13, we can apply Corollary 4.3, the first part gives Item 3.
+The second part, together with the fact that $G_{i}$ has property $(T)$, gives a short exact sequence
 $0\to H_{2}(G_{i};V)\to H_{2}(G_{i+1};V)\to V_{g_{i}}\to 0.$
@@
 We now prove A.
-The proof will be similar, but more technical, in that we need to leverage an additional prime, and the full power of Proposition 4.8, to make the group $G_{\alpha}$ simple.
+The proof will be similar, but more technical, in that we need to leverage an additional prime, and the full power of Proposition 4.13, to make the group $G_{x}$ simple.
 This improvement employs a trick from [OT13], where it was used to construct simple groups with large first $\ell^{2}$-Betti numbers.
-The following will be our starting point.
+The following will be our starting point, this is also the only place where we use Proposition 4.12 for the induction step.
 Lemma 6.2.
-Let $G$ be a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane. Let $q$ be a prime. Then there exists a quotient $G\to G_{1}$ with the following properties.
+Let $G$ be a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane.
+Let $q$ be a prime and let $\varepsilon>0$.
+Then there exists a quotient $G\to G_{0}$ with the following properties.
   1. 1.
-    $G_{1}$ is non-elementary hyperbolic with no non-trivial finite normal subgroups.
+    $G_{0}$ is non-elementary hyperbolic with no non-trivial finite normal subgroups.
   2. 2.
-    The quotient $G\to G_{1}$ induces isomorphisms in homology and cohomology with coefficients in $R[G_{1}]$-modules and degree at least $3$, where $R$ is any ring where $q$ is invertible.
+    The quotient $G\to G_{0}$ induces isomorphisms in homology and cohomology with coefficients in $R[G_{0}]$-modules and degree at least $3$, where $R$ is any ring where $q$ is invertible.
   3. 3.
-    $G_{1}$ is generated by elements $\{x_{1},\ldots,x_{m}\}$, each of which has order a power of $q$.
+    $G_{0}$ is generated by elements $\{a_{1},\ldots,a_{m}\}$, each of which has order a power of $q$.
   4. 4.
-    There is a finite set $\mathcal{E}$ of finite subgroups of $G_{1}$, such that for every unitary $G_{1}$-representation $V$, the second homology is
-    $H_{2}(G_{1};V)=\bigoplus_{E\in\mathcal{E}}V_{E}.$
+    There is a finite set $\mathcal{E}$ of finite subgroups of $G_{0}$, such that for every unitary $G_{0}$-representation $V$, the second homology is
+    $H_{2}(G_{0};V)=\bigoplus_{E\in\mathcal{E}}V_{E},$
+    and moreover
+    $\sum\limits_{E\in\mathcal{E}}|E|^{-1}<\varepsilon.$
 Proof.
-We start by choosing a generating set $\{x_{1}^{0},\ldots,x_{m}^{0}\}$ of $G$. For $i=0,\ldots,m$, we construct by induction a quotient $G\to G_{i/m}$ with the following properties.
+We start by choosing a generating set $\{a_{1}^{0},\ldots,a_{m}^{0}\}$ of $G$. For $i=1,\ldots,m$, we construct by induction a quotient $G\to G^{i}$ with the following properties.
   1. 1.
-    $G_{i/m}$ is non-elementary hyperbolic with no non-trivial finite normal subgroups.
+    $G^{i}$ is non-elementary hyperbolic with no non-trivial finite normal subgroups.
   2. 2.
-    The quotient $G\to G_{i/m}$ induces isomorphisms in homology and cohomology with coefficients in $R[G_{i/m}]$-modules and degree at least $3$, where $R$ is any ring where $q$ is invertible.
-In particular $G_{i/m}$ has cohomological dimension $16$ modulo $q$, and so it only contains $q$-torsion [Bie81, Proposition 4.11 on p.
+    The quotient $G\to G^{i}$ induces isomorphisms in homology and cohomology with coefficients in $R[G^{i}]$-modules and degree at least $3$, where $R$ is any ring where $q$ is invertible.
+In particular $G^{i}$ has cohomological dimension $16$ modulo $q$, and so it only contains $q$-torsion [Bie81, Proposition 4.11 on p.
 63].
   3. 3.
-    $G_{i/m}$ is generated by elements $\{x_{1}^{i},\ldots,x_{m}^{i}\}$, where $x_{1}^{i},\ldots,x_{i}^{i}$ have order a power of $q$.
+    $G^{i}$ is generated by elements $\{a_{1}^{i},\ldots,a_{m}^{i}\}$, where $a_{1}^{i},\ldots,a_{i}^{i}$ have order a power of $q$.
   4. 4.
-    There is a finite set of finite subgroups $\mathcal{E}_{i}$ of $G_{i/m}$, such that for every unitary $G_{i/m}$-representation $V$, the second homology is
-    $H_{2}(G_{i/m};V)=\bigoplus_{E\in\mathcal{E}_{i}}V_{E}.$
-Then the final group $G_{m/m}$ will be the group $G_{1}$ in the statement of the lemma, with $\mathcal{E}=\mathcal{E}_{m}$ and $x_{j}=x_{j}^{m}$.
+    There is a finite set of finite subgroups $\mathcal{E}_{i}$ of $G^{i}$, such that for every unitary $G^{i}$-representation $V$, the second homology is
+    $H_{2}(G^{i};V)=\bigoplus_{E\in\mathcal{E}_{i}}V_{E},$
+    and moreover
+    $\sum\limits_{E\in\mathcal{E}_{i}}|E|^{-1}<\varepsilon.$
+Then the final group $G^{m}$ will be the group $G_{0}$ in the statement of the lemma, with $\mathcal{E}=\mathcal{E}_{m}$ and $a_{j}=a_{j}^{m}$.
 The base case for $i=0$ is $G$ itself, with $\mathcal{E}=\emptyset$, where we do not need to check anything for the first three items, and Item 4 follows from Proposition 3.8 and Theorem 3.10.
-Suppose by induction that $G_{i/m}$ has been constructed.
-If $x_{i+1}^{i}$ is torsion, then it is $q$-torsion, so we set $G_{(i+1)/m}=G_{i/m}$, $x^{i+1}_{j}=x^{i}_{j}$ for all $j$, and $\mathcal{E}_{i+1}=\mathcal{E}_{i}$.
-Otherwise, consider the infinite virtually cyclic group $E_{G_{i/m}}(x_{i+1})$.
+Suppose by induction that $G^{i}$ has been constructed.
+If $a_{i+1}^{i}$ is torsion, then it is $q$-torsion, so we set $G^{i+1}=G^{i}$, $a^{i+1}_{j}=a^{i}_{j}$ for all $j$, and $\mathcal{E}_{i+1}=\mathcal{E}_{i}$.
+Otherwise, consider the infinite virtually cyclic group $E_{G^{i}}(a_{i+1})$.
 This has a unique maximal finite normal subgroup of order a power of $q$, and the quotient is isomorphic to $\mathbb{Z}$ (or possibly to $D_{\infty}$, in case $q=2$).
-We replace $x_{i+1}^{i}$ with a root $x^{i+1}_{i+1}$ that generates the $\mathbb{Z}$ quotient (or the index-$2$ subgroup of the $D_{\infty}$ quotient), so that $E_{G_{i/m}}(x^{i+1}_{i+1}):\langle x^{i+1}_{i+1}\rangle]$ is a power of $q$.
-We now let $G_{(i+1)/m}$ be the group-theoretic Dehn filling for the triple $(G_{i/m},E_{G_{i/m}}(x^{i+1}_{i+1}),\langle(x^{i+1}_{i+1})^{k}\rangle)$, where $k$ is a sufficiently large power of $q$.
-By Proposition 4.7, we see that $G_{(i+1)/m}$ satisfies Items 1 and 3, choosing $x^{i+1}_{j}=x^{i}_{j}$ for all $j\neq i+1$.
-We also choose $k$ to be large enough so that the orders of $x_{1}^{i},\ldots,x_{i}^{i}$ and their images $x_{1}^{i+1},\ldots,x_{i}^{i+1}$ have the same order, and each finite subgroup from $\mathcal{E}_{i}$ embeds into $G_{(i+1)/m}$.
+We replace $a_{i+1}^{i}$ with a root $a^{i+1}_{i+1}$ that generates the $\mathbb{Z}$ quotient (or the index-$2$ subgroup of the $D_{\infty}$ quotient), so that $E_{G^{i}}(a^{i+1}_{i+1}):\langle a^{i+1}_{i+1}\rangle]$ is a power of $q$.
+We now let $G^{i+1}$ be the group-theoretic Dehn filling for the triple $(G^{i},E_{G^{i}}(a^{i+1}_{i+1}),\langle(a^{i+1}_{i+1})^{k}\rangle)$, where $k$ is a sufficiently large power of $q$.
+By Proposition 4.12, we see that $G^{i+1}$ satisfies Items 1 and 3, choosing $a^{i+1}_{j}=a^{i}_{j}$ for all $j\neq i+1$.
+We also choose $k$ to be large enough so that the orders of $a_{1}^{i},\ldots,a_{i}^{i}$ and their images $a_{1}^{i+1},\ldots,a_{i}^{i+1}$ have the same order, and each finite subgroup from $\mathcal{E}_{i}$ embeds into $G^{i+1}$.
 Moreover, by Proposition 4.12 again, this triple is Cohen–Lyndon, so Corollary 4.3 applies, and proves Item 2.
-Finally, let $E\coloneqq E_{G_{i/m}}(x_{i+1}^{i+1})/\langle(x_{i+1}^{i+1})^{k}\rangle$.
-Corollary 4.3 and property $(T)$ for $G_{i/m}$ give a short exact sequence
-$0\to H_{2}(G_{i/m};V)\to H_{2}(G_{(i+1)/m};V)\to V_{E}\to 0,$
-so setting $\mathcal{E}_{i+1}=\mathcal{E}_{i}\cup\{E\}$, we obtain Item 4 and conclude. ∎
+Finally, let $E\coloneqq E_{G^{i}}(a_{i+1}^{i+1})/\langle(a_{i+1}^{i+1})^{k}\rangle$.
+Corollary 4.3 and property $(T)$ for $G^{i}$ give a short exact sequence
+$0\to H_{2}(G^{i};V)\to H_{2}(G^{i+1};V)\to V_{E}\to 0,$
+so setting $\mathcal{E}_{i+1}=\mathcal{E}_{i}\cup\{E\}$, we obtain the first part of Item 4. The second part follows again by induction, if we choose $k$ to be sufficiently large, and we conclude. ∎
 Proof of A.
-We proceed as in the proof of Theorem 6.1, using the prime $p\neq q$, but starting with the group $G_{1}$ from Lemma 6.2.
-The result will be a group $G_{\alpha}$ with the same properties as in Theorem 6.1, with the exception that cohomological dimension now has to be taken modulo $\{p,q\}$, and $b^{(2)}_{2}(G_{\alpha})$ is computed starting from
-$b^{(2)}_{2}(G_{1}\to G_{\alpha})=\dim_{\mathcal{N}(G_{\alpha})}\Bigl(\bigoplus_{E\in\mathcal{E}}\mathcal{N}(G_{\alpha})_{E}\Bigr)=\sum\limits_{E\in\mathcal{E}}|E|^{-1}.$
-Since this starting value is independent of $\alpha$, the same computations go through. For the finiteness of the $\ell^{2}$-Betti numbers of $G_{\alpha}$ we refer to Remark 6.3.
-In each induction step in the proof of Theorem 6.1, we chose arbitrary special elements $g_{i}\in G_{i}$.
+We proceed as in the proof of Theorem 6.1, using the prime $p\neq q$, but starting with the group $G_{0}$ from Lemma 6.2.
+The result will be a group $G_{x}$ with the same properties as in Theorem 6.1, with the exception that cohomological dimension now has to be taken modulo $\{p,q\}$, and $b^{(2)}_{2}(G_{x})$ is computed starting from
+$b^{(2)}_{2}(G_{0}\to G_{x})=\dim_{\mathcal{N}(G_{x})}\Bigl(\bigoplus_{E\in\mathcal{E}}\mathcal{N}(G_{\alpha})_{E}\Bigr)=\sum\limits_{E\in\mathcal{E}}|E|^{-1}.$
+We let this value be $\varepsilon<x$.
+The same computations go through, where now work with a base-$p$ representation of $x-\varepsilon$ instead.
+For the finiteness of the higher $\ell^{2}$-Betti numbers of $G_{x}$ we refer to Remark 6.3.
+In each induction step in the proof of Theorem 6.1, we chose special elements $g_{i}\in G_{i}$, using Proposition 4.13 with arbitrary $x,H$.
 We will now explain how to use the full power of Proposition 4.13 in the construction so that this time the colimit is simple.
-Enumerate the set $\mathcal{X}\coloneqq\{x_{1},\ldots,x_{m}\}\times G_{1}$.
-When the group $G_{i}$ has been built, let $(x,h)$ denote the $i$-th element of $\mathcal{X}$.
+Enumerate the set $\mathcal{A}\coloneqq\{a_{1},\ldots,a_{m}\}\times G_{0}$.
+When the group $G_{i}$ has been built, let $(a,h)$ denote the $i$-th element of $\mathcal{A}$.
 If $h$ is trivial, then we pick $g_{i}$ to be an arbitrary special element, and proceed as before.
-Otherwise, we use Proposition 4.8 to pick $g_{i}\in x^{-1}\left<\!\left<h\right>\!\right>_{G_{i}}$.
-We claim that now $x\in\left<\!\left<h\right>\!\right>_{G_{i+1}}\leqslant G_{i+1}$.
-Indeed, consider the image of $x$ in the quotient $G_{i+1}/\left<\!\left<h\right>\!\right>_{G_{i+1}}$.
-On the one hand, it has order a power of $q$, because $x$ has order a power of $q$ in $G_{1}$.
+Otherwise, we use Proposition 4.13 to pick $g_{i}\in a^{-1}\left<\!\left<h\right>\!\right>_{G_{i}}$, which is possible since $\left<\!\left<h\right>\!\right>_{G_{i}}$ is suitable by Lemma 4.10.
+We claim that now $a\in\left<\!\left<h\right>\!\right>_{G_{i+1}}\leqslant G_{i+1}$.
+Indeed, consider the image of $a$ in the quotient $G_{i+1}/\left<\!\left<h\right>\!\right>_{G_{i+1}}$.
+On the one hand, it has order a power of $q$, because $a$ has order a power of $q$ in $G_{0}$.
 On the other hand, it equals the image of $g_{i}$, which has order a power of $p$ in $G_{i+1}$, by construction of the quotient $G_{i}\to G_{i+1}$.
-Because $p\neq q$, we see that $x$ must be trivial in $G_{i+1}/\left<\!\left<h\right>\!\right>_{G_{i+1}}$, which proves our claim.
-As a result, for every $(x,h)\in\mathcal{X}$, either $h$ is trivial in $G_{\alpha}$, or $x$ lies in the normal closure of $h$ in $G_{\alpha}$.
-Because $x$ ranges over a generating set for $G_{1}$ and $h$ ranges over $G_{1}$, this shows that every non-trivial element of $G_{\alpha}$ normally generates, that is, $G_{\alpha}$ is simple.
+Because $p\neq q$, we see that $a$ must be trivial in $G_{i+1}/\left<\!\left<h\right>\!\right>_{G_{i+1}}$, which proves our claim.
+As a result, for every $(a,h)\in\mathcal{A}$, either $h$ is trivial in $G_{x}$, or $a$ lies in the normal closure of $h$ in $G_{x}$.
+Because $a$ ranges over a generating set for $G_{0}$ and $h$ ranges over $G_{0}$, this shows that every non-trivial element of $G_{x}$ normally generates, that is, $G_{x}$ is simple.
 ∎
@@
 Note that the constructions give some information about higher $\ell^{2}$-Betti numbers as well.
-Indeed, $G_{\alpha}$ is obtained from a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane, denoted $G$, by a sequence of group-theoretic Dehn fillings satisfying Corollary 4.3.
+Indeed, $G_{x}$ is obtained from a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane, denoted $G$, by a sequence of group-theoretic Dehn fillings satisfying Corollary 4.3.
 Using this and Proposition 2.2 we have
-$b^{(2)}_{n}(G_{\alpha})=b^{(2)}_{n}(G\to G_{\alpha})$
+$b^{(2)}_{n}(G_{x})=b^{(2)}_{n}(G\to G_{x})$
 for all $n\geqslant 3$.
-Since $G$ is of type $F$, this implies that all $\ell^{2}$-Betti numbers of $G_{\alpha}$ are finite.
-We know from Theorem 3.10 that $G$ has property $(T_{3})$, so Corollary 3.9 gives $b^{(2)}_{3}(G_{\alpha})=0$.
+Since $G$ is of type $F$, this implies that all $\ell^{2}$-Betti numbers of $G_{x}$ are finite.
+We know from Theorem 3.10 that $G$ has property $(T_{3})$, so Corollary 3.9 gives $b^{(2)}_{3}(G_{x})=0$.
 Moreover, $G$ satisfies the Singer conjecture [Lüc02, Corollary 5.16 on p.
 231], that is $b^{(2)}_{n}(G)=0$ for all $n\neq 8$ and $b^{(2)}_{8}(G)=\chi\geqslant 1$.
-This does not give information about $b^{(2)}_{n}(G\to G_{\alpha})$, however, assuming Lück’s approximation conjecture [Lüc02, Chapter 13], and repeating the construction ensuring that $G\to G_{\alpha}$ is injective on a large ball centred at the identity, we would have that $b^{(2)}_{n}(G\to G_{\alpha})$ is close to $b^{(2)}_{n}(G)$.
-It would then follow that $b^{(2)}_{n}(G_{\alpha})$ is close to 0, for $n\geqslant 4,n\neq 8$, and close to $\chi$ for $n=8$.
+This does not give information about $b^{(2)}_{n}(G\to G_{x})$, however, assuming Lück’s approximation conjecture [Lüc02, Chapter 13], and repeating the construction ensuring that $G\to G_{x}$ is injective on a large ball centred at the identity, we would have that $b^{(2)}_{n}(G\to G_{x})$ is close to $b^{(2)}_{n}(G)$.
+It would then follow that $b^{(2)}_{n}(G_{x})$ is close to 0, for $n\geqslant 4,n\neq 8$, and close to $\chi$ for $n=8$.
 Remark 6.4.
-Taking $V=\mathbb{C}$ to be the trivial unitary representation, the proofs of Theorems 6.1 and A, and Proposition 2.2, show that $H_{2}(G_{\alpha};\mathbb{C})$ is infinite-dimensional.
+Taking $V=\mathbb{C}$ to be the trivial unitary representation, the proofs of Theorems 6.1 and A, and Proposition 2.2, show that $H_{2}(G_{x};\mathbb{C})$ is infinite-dimensional.
 It follows that the groups $G_{\alpha}$ do not have type $FP_{2}(\mathbb{Q})$.
 7 An elementary construction
-We conclude with the elementary construction giving D. We start with two basic product formulas.
+We conclude with the elementary construction giving F. We start with two basic product formulas.
 Proposition 7.1 ([Lüc02, Theorem 6.54(5) on p. 266]).
@@
 ∎
-Proof of D.
+Proof of F.
 Let $C_{x}$ be the groups from Proposition 7.3.
@@
 $H_{n}((R_{x}\times F_{2}^{3});\mathcal{N}(G_{x}))^{2}\to H_{n}(G_{x};\mathcal{N}(G_{x}))\to H_{n-1}(C_{x};\mathcal{N}(G_{x}))\to H_{n-1}(R_{x}\times F_{2}^{3});\mathcal{N}(G_{x}))^{2}.$
-Taking $\dim_{G_{x}}$ and $n\leqslant 3$, the first and last term give $2b^{(2)}_{n}(R_{x}\times F_{2}^{3})=2b^{(2)}_{n-1}(R_{x}\times F_{2}^{3})=0$, by Proposition 3.11, so
+Taking $\dim_{\mathcal{N}(G_{x})}$ and $n\leqslant 3$, the first and last term give $2b^{(2)}_{n}(R_{x}\times F_{2}^{3})=2b^{(2)}_{n-1}(R_{x}\times F_{2}^{3})=0$, by Proposition 3.11, so
 $b^{(2)}_{n}(G_{x})=b^{(2)}_{n-1}(C_{x}).$
@@
 Corollary 7.5.
-Let $(G_{x})_{x\in\mathbb{R}_{>0}}$ be as in D. Then the groups $G_{x}\ast\mathbb{Z}$ are pairwise non-measure equivalent.
+Let $(G_{x})_{x\in\mathbb{R}_{>0}}$ be as in F. Then the groups $G_{x}\ast\mathbb{Z}$ are pairwise non-measure equivalent.
 Proof.